### v1

# Noise-based reward-modulated learning

Abstract
Recent advances in reinforcement learning (RL) have led to significant improvements in task performance. However, training neural networks in an RL regime is typically achieved in combination with backpropagation, limiting their applicability in resource-constrained environments or when using non-differentiable neural networks. While noise-based alternatives like reward-modulated Hebbian learning (RMHL) have been proposed, their performance has remained limited, especially in scenarios with delayed rewards, which require retrospective credit assignment over time. Here, we derive a novel noise-based learning rule that addresses these challenges. Our approach combines directional derivative theory with Hebbian-like updates to enable efficient, gradient-free learning in RL. It features stochastic noisy neurons which can approximate gradients, and produces local synaptic updates modulated by a global reward signal. Drawing on concepts from neuroscience, our method uses reward prediction error as its optimization target to generate increasingly advantageous behavior, and incorporates an eligibility trace to facilitate temporal credit assignment in environments with delayed rewards. Its formulation relies on local information alone, making it compatible with implementations in neuromorphic hardware. Experimental validation shows that our approach significantly outperforms RMHL and is competitive with BP-based baselines, highlighting the promise of noise-based, biologically inspired learning for low-power and real-time applications.
1 Introduction
Noise-based learning methods, which leverage stochastic fluctuations in neural activity [15, 20, 54, 52, 13] or synaptic weights [10, 57, 5, 8] to approximate gradients, offer a pathway toward simpler and more light-weight optimization techniques, that may reflect mechanisms within the brain itself [43, 29]. Given the prevalence of noise in biological neural systems [9, 45, 31], using noise as a mechanism for learning synaptic weights is an area of growing interest, with reward-modulated Hebbian learning (RMHL) [26, 32] as a promising candidate. RMHL, a form of three-factor Hebbian learning [17, 24], offers a potentially powerful mechanism for credit assignment without explicit backpropagation of exact gradients.
Despite theoretical advances, few noise-based methods have been adapted to real-world tasks. Some attempts which have been made include application to control problems [4], for spatiotemporal pattern generation [23], and for training non-differentiable spiking neural networks training [12], and these particularly struggle with delayed reward regimes. This gap has limited the integration of noise-based approaches for effective credit assignment in temporally extended setups.
In this work, we address this gap by proposing a novel noise-driven learning mechanism for reinforcement learning (RL), referred to as noise-based reward-modulated learning (NRL), which leverages noise to approximate gradients for synaptic updates derived from reward maximization. NRL integrates key concepts from neuroscience, such as eligibility traces [19], which link past neural states to future rewards, and reward prediction error (RPE) [41, 40, 34], which drives learning toward increasingly more advantageous behaviors. We employ directional derivatives for gradient approximation, which is realized through stochastic neurons that are compatible with the inherently noisy nature of brains and other physical systems. NRL estimates the directional derivative via two forward passes: a “noisy” pass with stochastic neurons and a “clean” pass without noise. In scenarios where a noiseless pass is infeasible, as in noisy analog hardware, multiple noisy passes can be averaged to approximate the clean pass, maintaining performance. We further validate our approach on RL benchmarks with both immediate- and delayed-reward problems. Finally, we discuss the broader implications for machine learning, neuromorphic computing, and neuroscience.
2 Methods
2.1 Agent-environment setup and learning signals
We consider an agent interacting continuously with a dynamic environment, whose state at time $t$ is denoted by $S_{t}$. The agent observes $S_{t}$ and responds by selecting actions, $A_{t}$, which influence the environment. Each action or sequence of actions yields feedback in the form of positive or negative rewards $r_{t}$. This observe-act loop provides a feedback loop which can be leveraged to enable an agent to learn and adapt to the changing environment, aiming to maximize the rewards associated with its actions.
The agent’s decision-making is guided by a policy $\pi$, representing the probability of choosing an action, $a$, in the state $S_{t}$, i.e., $\pi(a\mid S_{t})$. This policy reflects the agent’s understanding of optimal actions in each state. To determine an action, the agent computes a probability distribution over the set of possible actions, modeled as a categorical distribution, where $\sum_{a}\pi(a\mid S_{t})=1$, such that
$P(A_{t}=a\mid S_{t})=\pi(a\mid S_{t})\,.$ (1)
Rewards, $r_{t}$, from the environment, received after each action or sequence of actions, are used to compute the learning signal. Internally, the agent maintains a prediction of reward, denoted by $\bar{r}_{t}$. This prediction could be formed by an arbitrary computation, but for simplicity it is modeled as a running average over recent rewards
$\bar{r}_{t}=\bar{r}_{t-1}+\lambda\left(r_{t}-\bar{r}_{t-1}\right)$ (2)
where $\lambda$ is a smoothing factor. The mismatch between the predicted reward and the actual reward, denoted by
$\delta r_{t}=r_{t}-\bar{r}_{t}$ (3)
serves as the primary learning signal. It can be interpreted as a reward prediction error (RPE) in neuroscience contexts and can be thought of as a global dopaminergic signal that modulates learning in the brain [41, 40, 34]. While conceptually similar to biological RPEs, this formulation is a simplification due to the straightforward running average used for reward prediction. Nonetheless, it effectively captures the core ideas of the proposed learning mechanism. In the Discussion, we explore how this prediction error could be refined.
2.2 Derivation of the learning rule
This section presents a complete derivation of the NRL update rule. We begin by establishing a gradient-based learning rule rooted in an optimization target that aims to maximize unexpected rewards from the environment by increasing the reward prediction error. Maximizing the RPE rather than the direct reward enables the agent to adapt dynamically to changing environments, improve long-term performance by continuously seeking rewards that exceed expectations, and maintain robustness against variations in reward structures [48, 36]. Building on this foundation, we transition to a gradient-free formulation by leveraging a directional derivative framework, where stochastic noise within the network approximates gradients. This eliminates the need for explicit backpropagation and feedback phases, enabling learning through forward passes alone. Finally, we extend our noise-based approach to handle scenarios with delayed rewards, a hallmark of real-world problems. This extension enables NRL to adapt to environments where feedback is only obtained after a series of actions.
2.2.1 Gradient-based learning rule
The agent is represented by a neural network which processes input observations and generates a probability distribution over possible actions. This neural network consists of multiple layers, each performing a linear transformation of its inputs, followed by a non-linear transformation as
$x^{l}=f(h^{l})=f(W^{l}x^{l-1})$ (4)
where $x^{l}$ represents the output of layer $l$, $h^{l}$ represents the layer pre-activation, $W^{l}$ is the weight matrix, and $f(\cdot)$ is a non-linear activation function.
In this setup, the goal is to optimize the weights $W=\{W^{1},\ldots,W^{l}\}$ across the $L$ layers in the neural network to maximize the received rewards. Our derivation begins by introducing a generic variable $\theta$, which will later be mapped to the specific neural network parameters $W$. The primary learning objective is to maximize the expected RPE, $\delta r_{t}$, which aims to achieve higher-than-anticipated rewards. For now, let us consider learning to maximise RPE at a specific timepoint $t$. Since the reward prediction is a running average of recent rewards, the learning rule seeks to outperform previously received rewards. We express this objective in terms of the policy as
$J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\delta r_{t}\right]\,.$ (5)
To optimize $J(\theta)$, we can incrementally update the weights using the gradients with respect to $\theta$ as
$\theta\leftarrow\theta+\eta\nabla_{\theta}J(\theta)$ (6)
where $\eta$ is the learning rate, controlling the size of each update step.
Applying the policy gradient theorem [49], which utilizes the likelihood-ratio method, we express the gradient of the objective as
$\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\nabla_{\theta}\log\pi_{\theta}\left(a\mid S_{t}\right)\delta r_{t}\right]\,.$ (7)
For empirical estimation, in the case of a single sample, we approximate the gradient as
$\nabla_{\theta}J(\theta)\approx\nabla_{\theta}\log\pi_{\theta}\left(a\mid S_{t}\right)\delta r_{t}\,.$ (8)
With this approximation, we define the parameter updates as
$\theta\leftarrow\theta+\eta\;\nabla_{\theta}\log\pi_{\theta}\left(a\mid S_{t}\right)\;\delta r_{t}.$ (9)
Equation (9) resembles the REINFORCE update rule [54] but differs by using the reward prediction error (RPE), $\delta_{r}$, as the learning signal instead of cumulative rewards over full trajectories. This RPE-based approach leverages immediate feedback from rewards as they are obtained rather than requiring a full trial completion to estimate the policy gradient. It shares conceptual similarities with actor-critic methods [1], where the policy is adjusted using a temporal difference error. However, we approximate future rewards with a running average of past rewards instead of a critic network, maintaining adaptive feedback benefits with a simpler implementation.
2.2.2 Noise-based learning rule
The learning rule derived in the previous section still relies on gradient descent. To avoid using backpropagation (BP) to compute the gradients, we propose a noise-based alternative that extends Equation (9). This approach leverages gradient approximation via directional derivatives, as detailed in Appendix A, enabling a theoretically rigorous derivation of noise-driven learning.
A directional derivative quantifies the rate of change of a function in a specified direction. In neural networks, we implement this concept by introducing random noise into parameters. By comparing the network’s parameters with and without this noise, we obtain an estimate of the gradient direction. To formalize this, we express the gradient term in terms of directional derivatives using Theorem 31 in Appendix A as
$\nabla p(\theta)$ $=\nabla_{\theta}\log\pi_{\theta}(a\mid S_{t})=n\mathbb{E}\left[v\nabla_{v}p(\theta)\right]$ (10)
where $v=\ifrac{\epsilon}{||\epsilon||}$ is a normalized direction vector derived from noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$ with $n$ the number of parameters.
We may expand the above via Theorem A.4, and add a difference term (parameters with noise versus without noise) to approximate the gradient by sampling under an empirical distribution
$\nabla p(\theta)=n\sum_{i=1}^{K}\left[\frac{\epsilon^{(i)}}{||\epsilon^{(i)}||^{2}}\left(p\bigl{(}\tilde{\theta}^{(i)}\bigr{)}-p\bigl{(}\theta\bigr{)}\right)\right]\,.$ (11)
where $K$ denotes the number of samples and $\tilde{\theta}^{(i)}=\theta+\epsilon^{(i)}$ the noise-perturbed parameters. In practice, we consider $K=1$, analogous to single-sample updates in stochastic gradient descent. Here, $p\bigl{(}\tilde{\theta}\bigr{)}=\log\pi_{\tilde{\theta}}(a\mid S_{t})$ represents the log of the noise-perturbed output, while, $p\left(\theta\right)=\log\pi_{\theta}(a\mid S_{t})$ is the log of the noise-free output. Thus, $\rho=\log\pi_{\tilde{\theta}}(a\mid S_{t})-\log\pi_{\theta}(a\mid S_{t})$ captures the impact of the noise on the policy. Putting this together for the $K=1$ case, we obtain our noise-based learning rule
$\theta\leftarrow\theta+\eta\delta_{r}\epsilon^{*}\rho$ (12)
with $\epsilon^{*}=\ifrac{\epsilon}{||\epsilon||^{2}}$. For convenience, we absorbed the term $n$ into the learning rate $\eta$.
We can extend this by introducing noise into the neurons instead of the weights of the network, allowing localized perturbations within network layers. Given a layer transformation as in Equation (4), adding noise at the neuron level is represented as
$\tilde{x}^{l}=f\left(\tilde{h}^{l}+\xi^{l}\right)=f\left(W^{l}\tilde{x}^{l-1}+\xi^{l}\right)$ (13)
where $\xi^{l}\sim\mathcal{N}(0,\beta\sigma^{2}I_{m^{l}})$ is the injected noise, $\beta$ is some arbitrarily small noise scale, and $m^{l}$ is the number of neurons in layer $l$. We use the notation $\tilde{x}^{l}$ and $\tilde{h}^{l}$ to denote perturbed inputs and pre-activations, respectively, which may also result from perturbations of previous network layers on which layer $l$ depends.
To formulate the learning rule for specific layer parameters $W^{l}\in W$, we re-express the gradient in terms of the pre-activations instead of the parameters themselves. To do so, we first rewrite the gradient term as
$\nabla p\left(W^{l}\right)=\nabla_{W^{l}}\log\pi_{\theta}\left(a\mid S_{t}\right)$ (14)
and apply the chain rule of differentiation to break down the gradient with respect to $W^{l}$ as
$\nabla p\left(W^{l}\right)=\frac{\partial p}{\partial\tilde{h}^{l}}\frac{\partial\tilde{h}^{l}}{\partial W^{l}}=\nabla p\left(\tilde{h}^{l}\right)\tilde{x}^{l-1}\,.$ (15)
Building on the theory developed above, we derive the gradient $\nabla p(\tilde{h})$, leading to the layer-specific weight update rule
$W^{l}\leftarrow W^{l}+\eta\delta r_{t}\bar{\xi}^{l}{\rho}\tilde{x}^{l-1}$ (16)
with ${\rho}=\log\pi_{\tilde{W}}(a\mid S_{t})-\log\pi_{W}(a\mid S_{t})$ and $\bar{\xi}^{l}=\ifrac{\xi^{l}}{||\xi^{l}||^{2}}$. Here, $\pi_{\tilde{W}}(a\mid S_{t})$ represents the network’s output when noise is injected into the neurons, while $\pi_{W}(a\mid S_{t})$ corresponds to the output of the noiseless network.
2.2.3 Learning from sparse rewards
In our derivation above, synaptic updates occur at every time step, implying that learning signals and rewards are received continuously. However, rewards are usually provided after completing a sequence of actions or upon reaching specific milestones, leading to more sparsely spaced feedback. To handle this scenario, we modify our learning rule to accommodate sparse rewards arriving at arbitrary times.
Starting from Equation (9), we redefine the synaptic weight update to include time explicitly
$W^{l}_{t+1}=W^{l}_{t}+\eta\;\delta r_{t}\;\bar{\xi}^{l}_{t}p_{t}\tilde{x}^{l-1}_{t}$ (17)
where $p_{t}=\log\pi_{\tilde{W}}(a\mid S_{t})-\log\pi_{W}(a\mid S_{t})$. Next, we assume sparse rewards arriving at arbitrary times $\tau$, and thus synaptic updates only occur at this moment, denoted by $W^{l}_{\tau}$. The learning rule now accumulates information since the last reward was received, updating weights by integrating this information over time
$W^{l}_{\tau}\leftarrow W^{l}_{\tau_{0}}+\eta\;\delta r_{\tau}\;\sum_{t=\tau_{0}}^{\tau}\bar{\xi}^{l}_{t}{p}_{t}\tilde{x}^{l-1}_{t}$ (18)
where $\tau_{0}$ represents the time when the previous reward was obtained.
We further represent this accumulated sum using an eligibility trace, which captures the history of neural activity and noise contributions since the last reward, defined as
$e_{\tau}=\sum_{t=\tau_{0}}^{\tau}\bar{\xi}^{l}_{t}{p}_{t}\tilde{x}^{l-1}_{t}\,.$ (19)
such that (18) may be expressed as
$W^{l}_{\tau}\leftarrow W^{l}_{\tau_{0}}+\eta\;\delta r_{\tau}\;e_{\tau}\,.$ (20)
Thus, our final learning rule consists of two primary components: (i) an eligibility trace, defined in Equation (19), which accumulates local information over time at each time step, and (ii) a synaptic update, defined in Equation (20), triggered upon reward receipt, which modulates the eligibility trace to adjust the synaptic weights.
Eligibility traces act as a mechanism to connect past actions with future rewards [19]. They capture neural activity and other local variables over time, signaling potential synaptic changes. Upon receiving a reward, these traces are modulated by the reinforcement signal, resulting in synaptic updates. Some models view eligibility traces as decaying cumulative activity [21, 48]), while others treat them as a full activity history [32], which aligns with our formulation. The eligibility trace efficiently tracks neural information between rewards, facilitating the assignment of credit to past actions.
2.3 Neural network setup
At the core of the agent’s learning and decision-making process, we use a feedforward neural network with $L$ layers, where the transformation applied to each hidden layer is given by
$\tilde{x}^{l}$ $=f\left(W^{l}\tilde{x}^{l-1}+\xi^{l}\right)$ (21)
and the readout in the output layer is given by
$\tilde{y}$ $=s\left(W^{l}\tilde{x}^{l-1}+\xi^{l}\right)$ (22)
where $a^{l}$ and $x^{l}$ denote the layer pre-activation and output (with $x_{0}$ as the network input), $y$ is the network output, $\xi^{l}\sim\mathcal{N}(0,\beta\sigma^{2}I_{m^{l}})$ is noise added to the neurons, with $m^{l}$ representing the number of units in layer $l$. This is referred to as a “noisy pass”. The function $f(\cdot)$ represents a non-linear activation function applied to each layer. In our case, we use the LeakyReLU activation function such that $f(x)=x$ if $x\geq 0$ and $f(x)=\alpha x$ with $0<\alpha\ll 1$, otherwise. The function $s(\cdot)$ represents a softmax transformation ${s}_{i}(x)=\ifrac{e^{x_{i}}}{\sum_{j=1}^{m^{l}}e^{x_{j}}}$, converting the network’s output into action probabilities. We also define a “clean pass,” where no noise is present in the network, as
$x^{l}$ $=f\left(W^{l}x^{l-1}\right)$ (23)
$y$ $=s\left(W^{l}x^{l-1}\right)$ (24)
2.4 Experimental validation
We validate our approach across various environments and compare its performance with established baselines. Each environment presents an episodic problem, where the agent’s learning is guided solely by positive or negative rewards. We test in both instantaneous and delayed reward settings. In the delayed reward setting, a single reward is given after a sequence of actions, challenging the agent to assign credit retrospectively. For the instantaneous reward setting, we use the Reaching problem [18], while the delayed reward setting uses the Cartpole [1] and Acrobot [47] problems. We use implementations given by the libraries OpenAI Gym [3] and NeuroGym [33] for the different environments.
We compare NRL to two baselines: an exact-gradient version of NRL (similar to an actor-only variant of actor-critic which relies on a running average of past rewards as the prediction error), which serves as an ”optimal performance” benchmark and an RMHL approach. The first baseline uses BP for gradient computation, updating eligibility traces as
$e_{\tau}=\sum_{t=\tau_{0}}^{\tau}\nabla_{W}\log\pi_{W}\left(a\mid S_{t}\right)\,$ (25)
The second baseline, motivated by the noise-based nature of NRL, is the RMHL rule from [26] with an explicit-noise approximation – the original version where noise is inferred from neural activities proved too unstable for the problems considered here. For delayed rewards, we adapt it similarly to [32], updating eligibility traces as
$e_{\tau}=\sum_{t=\tau_{0}}^{\tau}\xi^{l}_{t}\tilde{x}^{l-1}_{t}\,.$ (26)
In environments with high reward variability, like Cartpole and Acrobot, we stabilize synaptic updates by scaling the RPE by dividing by $r_{\tau}$. Hyperparameter values and training details are provided in Appendix B. All experiments are implemented in Python using PyTorch [37]. Our models and scripts are available for reproducibility at https://github.com/jesusgf96/noise-based-reward-modulated-learning.
3 Results
3.1 Solving instantaneous and delayed reward problems
To ensure comparability with RMHL methods, which typically utilize single-hidden-layer networks, we first conduct experiments with a one-hidden-layer neural network for BP, NRL, and RMHL. Training details are provided in Appendix B. In Section 3.2, we extend the comparison to deeper networks with multiple hidden layers.
First, the Reaching problem [18], visualized in Figure 1C, is an instantaneous reward problem that requires an agent to reach and stay at a target on a 1D ring by moving left, right, or staying still. At each step, the agent observes both the target’s position and its own, and it is rewarded based on its proximity to the target over fixed-duration trials. Average across-trial and final performance (mean of the last 50 trials) are shown in Figures 1A and 1B, respectively.
Figure 1: Performance on benchmarks. A, B, C: Reaching problem. D, E, F: Acrobot problem. G, H, I: Cartpole problem. Left panels: Performance across trials averaged over 5 runs. Centre panels: Final performance (mean of the last 50 trials), averaged over 5 runs. Right panels: Problem visualization.
Second, the Acrobot problem [47], visualized in Figure 1F, involves delayed reward and requires controlling a two-link robotic arm to reach a target height. At each time step, the agent observes the angles and angular velocities of the two links and chooses one action: clockwise torque, counterclockwise torque, or no torque. Rewards are based on the speed of completion, with a maximum time. Average across-trial and final performance (mean of the last 50 trials) are shown in Figures 1D and 1E, respectively.
The third and most challenging problem, the Cartpole problem [1], visualized in Figure 1I, is a delayed reward problem where an agent must balance a pole on a cart by moving left or right. At each time step, the agent observes the cart’s position and velocity, along with the pole’s angle and angular velocity, and acts accordingly. Performance is measured by the time the pole remains balanced, with a maximum time. Average across-trial and final performance (mean of the last 50 trials) are shown in Figures 1G and 1H, respectively.
For all three tasks, we can observe that NRL reaches the performance level of BP whereas RMHL fails to do so.
3.2 Scalability to deeper networks
Figure 2: Performance on deeper networks – Acrobot. A, B: 2-hidden layer networks. C, D: 3-hidden layer networks. Left panels: Performance across trials averaged over 5 runs. Right panels: Final performance (mean of the last 50 trials), averaged over 5 runs.
Figure 3: Performance on deeper networks – Cartpole. A, B: 2-hidden layer networks. C, D: 3-hidden layer networks. Left panels: Performance across trials averaged over 5 runs. Right panels: Final performance (mean of the last 50 trials), averaged over 5 runs.
Here, we demonstrate that NRL can effectively assign credit in neural networks with multiple hidden layers – a challenging scenario where most biologically plausible algorithms struggle. We use both the Acrobot and the Cartpole tasks for this purpose, as they present a delayed reward problem, making them a more realistic test for credit assignment in reinforcement learning. In this comparison, we include the same baselines, BP and RMHL. Figure 2 displays the results for the Acrobot task and Figure 3 displays the results for the Cartpole tasks. The left panels of these two figures show performance across trials, averaged over 5 runs. In contrast, the right panels display the final performance (mean of the last 50 trials), also averaged over 5 runs. The results for neural networks with two hidden layers are shown in Figures 2A and B, and 3A and B, and those for networks with three hidden layers are presented in Figures 2C and D, and 3C and D.
Our results indicate that NRL successfully learns to solve the tasks, achieving performance comparable to BP. In contrast, RMHL struggles with credit assignment in deeper networks. However, NRL requires more trials to converge as the network depth increases—an expected outcome due to the stochastic nature of the updates [20]. A similar trend is observed with BP, though to a lesser extent, as its gradient-based updates inherently provide more directed adjustments.
3.3 Using only noisy passes
In NRL, ${p}_{t}$ is calculated as the difference in the network’s output between the clean and noisy passes. However, a completely noiseless network is not feasible in more biophysically realistic scenarios. Instead, the clean network’s output can be approximated by averaging the outputs from multiple noisy passes, as the injected noise averages to zero over time $\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\xi_{i}^{l}=0$. This assumes that the network dynamics are faster than the environment dynamics, allowing the network to perform multiple forward passes before the environment changes.
Figure 4: Learning using only noisy passes – Acrobot. A: Clean pass approximation error. Each data point is computed with an absolute error, averaged over 500 timesteps using the Acrobot problem. B: Acrobot problem. Performance across trials, averaged over 5 runs, with mean, minimum, and maximum values displayed.
To evaluate the accuracy of this approximation, we employ a single hidden-layer network and the Acrobot problem. The difference between the clean pass output and the averaged noisy passes output was calculated for each timestep and averaged over 500 timesteps, as shown in Figure 4A. Furthermore, Figure 4B illustrates the performance on the Acrobot problem using only noisy forward passes, starting with the minimum of 2 noisy passes. We also extended this evaluation to 10 noisy passes to explore the impact of increasing the number of passes, showing slightly faster initial convergence.
4 Discussion
In this work, we introduced NRL as a novel noise-driven learning mechanism for neural networks that bridges RL principles with Hebbian-like updates, providing a gradient-free alternative to traditional approaches. By leveraging directional derivatives and stochastic noisy neurons to approximate gradients [44, 53, 52], our method produces local synaptic updates that are modulated by a global reinforcement signal. This noise-based rule offers an efficient approach for neural adaptation, particularly suited to real-time and resource-limited environments.
NRL shares similarities with node perturbation methods [15, 20, 54, 52, 13], which inject noise into neurons to adjust weights based on state and performance fluctuations. However, unlike node perturbation –which requires constant feedback or reward for learning– NRL is well-suited for delayed-reward setups, a common scenario in real-world applications. By leveraging eligibility traces [19], NRL can effectively assign credit to past neural states even when actions and outcomes are separated in time.
Furthermore, NRL shares conceptual similarities with forward gradient methods [2], which also estimate gradients using directional derivatives, stochastic perturbations, and forward-mode differentiation. Both frameworks leverage noise-driven updates, focusing on solving real-world tasks. However, while forward gradient methods rely solely on random perturbations to approximate gradients, NRL refines updates by incorporating reinforcement learning principles, using RPEs to directly modulate synaptic weights. Recent advances in forward gradient techniques, such as employing local auxiliary networks to produce informed gradient guesses [16], have demonstrated substantial improvements in alignment with exact gradients, narrowing the performance gap with backpropagation. Similarly, these variance-reduction strategies could inspire further refinements in our approach, enhancing the quality of noise-based updates and improving scalability to more complex architectures.
Another increasingly popular approach for gradient estimation during optimization is zeroth-order (ZO) optimization [28, 6]. Like NRL, common ZO optimization relies on perturbation-driven strategies, utilizing differences in function values resulting from input perturbations to approximate gradients. However, as with forward gradient methods, NRL distinguishes itself by incorporating elements, such as RPEs and eligibility traces, to learn in delayed reward scenarios. Unlike forward gradient methods, which rely on directional derivatives and access to the model’s structure, ZO methods estimate gradients purely from function evaluations, without requiring any knowledge of the network’s internal structure or differentiability. This makes ZO methods particularly well-suited for black-box optimization tasks, a flexibility that our approach also shares.
The results demonstrate the effectiveness of NRL, significantly outperforming the comparable RMHL approach [26, 32] in both instantaneous and delayed reward tasks. While RMHL methods aim to align neural dynamics with observed brain dynamics rather than optimizing learning, our learning rule is rigorously derived to perform gradient approximation. As a result, our learning rule introduces two key innovations. First, normalizing the noise, $\bar{\xi}^{l}_{t}=\ifrac{\xi^{l}_{t}}{||\xi^{l}_{t}||^{2}}$, allows for smaller noise magnitudes that minimally impact agent behavior while still contributing effectively to the learning process. Second, an additional term, ${p}_{t}=\left[\log\pi_{\tilde{\theta}}(a|S_{t})-\log\pi_{\theta}(a|S_{t})\right]$, captures the influence of noise on the network output, thereby enhancing credit assignment accuracy.
Our results further demonstrate that NRL effectively assigns credit even in deep networks with multiple hidden layers, a challenge for many biologically plausible algorithms. As network depth increases, NRL’s performance advantage over RMHL becomes more pronounced. In fact, RMHL struggles with credit assignment in networks with more than one hidden layer, as evidenced by its minimal performance improvements during training. Nevertheless, NRL does require more trials to converge in deeper networks, which is an expected consequence of its stochastic update mechanism [20].
Compared to traditional BP-based methods, NRL approximates gradients rather than relying on exact gradients [27, 51, 39, 38], offering a significant advantage in computationally constrained scenarios—such or when using non-differentiable networks like spiking neural networks [55, 50]. The local-update structure of NRL provides compatibility with neuromorphic hardware [42], where efficient, low-power computations are essential. Neuromorphic systems can also inherently simulate noisy neural activity, enabling energy-efficient learning in applications where real-time adaptation is key.
The inherent noise in NRL is also advantageous for RL, as it naturally facilitates exploration-exploitation balance [22, 25]. Injected noise perturbs neural states, promoting exploratory deviations from expected values and thus introducing variations in the agent’s policy.
Our work further draws inspiration from neuroscience and underscores the importance of biologically plausible approaches [56, 30] for advancing learning algorithms. The reliance on RPEs as learning signals emulates the dopaminergic reward signals observed in the brain [41, 40], crucial for synaptic plasticity. The noise-based rule, coupled with eligibility traces, also aligns with biological processes that enable learning under uncertainty [14] and reflect the noisy nature of the brain [9, 45, 31]. Notice that the definition of RPEs used in this work is a simplification, relying on a straightforward running average for reward prediction. While this is sufficient to capture the core ideas of the proposed learning mechanism, more sophisticated approaches could further enhance reward estimation. One such extension is Temporal Difference (TD) learning [46, 48], which provides a more principled method for predicting expected future rewards and aligns more closely with biological dopaminergic activity [35]. A natural implementation would be an actor-critic architecture, where a critic network estimates the value function (e.g., expected future rewards), refining the reward prediction error (RPE).
One limitation of NRL is its dependence on multiple forward passes per synaptic update, which, though parallelizable, reduces biological plausibility. Future work could explore efficient ways to estimate noise impacts without explicitly needing separate clean and noisy passes. Additionally, while BP may offer faster convergence in unrestricted computational settings, recent research, including this study and others [7, 11], show promise in bridging this gap for specific networks and tasks.
Our findings suggest that noise-based learning strategies could drive advances in machine learning by enabling efficient, local, and energy-conscious algorithms, particularly suited for neuromorphic hardware and real-time applications. This research also offers valuable insights into how learning might occur in the brain, enriching our understanding of both artificial and biological intelligence.
Acknowledgments
This publication is part of the DBI2 project (024.005.022, Gravitation), which is financed by the Dutch Ministry of Education (OCW) via the Dutch Research Council (NWO).
Appendix A Gradient approximation using directional derivatives
Consider a function $f(\theta)$, where $\theta=(W_{1},\ldots,W_{L})$ are its parameters. The gradient of this function is defined as follows:
Definition A.1.
The gradient $\nabla f$ is a vector indicating the direction of the steepest ascent of the function $f$, with components as partial derivatives of $f(\theta)$:
$\nabla f=\left[\frac{\partial f}{\partial\theta}\right]^{\top}=\left[\frac{\partial f}{\partial W_{1}},\ldots,\frac{\partial f}{\partial W_{L}}\right]^{\top}\,.$ (27)
While $\nabla f$ captures the rate of change of $f$ in the steepest direction, a directional derivative gives the rate of change in a specified direction. For a unit vector $v=\ifrac{\epsilon}{||\epsilon||}$, normalized via the Euclidean norm, we define the directional derivative as:
Definition A.2.
The directional derivative of $f(\theta)$ along a unit vector $v=(v_{1},\ldots,v_{n})$ is defined by the limit
$\nabla_{v}{f}(\theta)=\lim_{h\to 0}{\frac{f(\theta+hv)-f(\theta)}{h}}\,,$ (28)
where $h$ is a small step size.
For a sufficiently small $h$ we approximate the directional derivative as
$\nabla_{{v}}{f}(\theta)=\frac{f(\theta+h\epsilon)-f(\theta)}{h||\epsilon||}\,.$ (29)
This directional derivative is essentially a projection of $\nabla f$ in the direction $v$, following the relation:
$\nabla_{v}{f}(\theta)=v\cdot\nabla{f}(\theta).$ (30)
We can now formally demonstrate how gradients can be approximated using directional derivatives.
Theorem A.3.
Let $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$, drawn from some probability distribution $p(\epsilon)$, and $n$ the number of dimensions in $\theta$. Exact gradients can be written in terms of directional derivatives using expectations
$\nabla f(\theta)=n\mathbb{E}_{p(\epsilon)}\left[v\nabla_{v}f(\theta)^{\top}\right]\,.$ (31)
Proof.
$\nabla f(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[v\nabla_{v}f(\theta)^{\top}\right]\,$
$=n\mathbb{E}_{p(\epsilon)}\left[vv\nabla f(\theta)^{\top}\right]\,$
$=n\mathbb{E}_{p(\epsilon)}\left[vv\right]\nabla f(\theta)^{\top}\,$
$=n\mathbb{E}_{p(\epsilon)}\left[\frac{\epsilon}{||\epsilon||}\frac{\epsilon^{\top}}{||\epsilon||}\right]\nabla f(\theta)\,$
$=n\frac{1}{n}I_{n}\nabla f(\theta)=\nabla f(\theta)\,$
as we assume $p(\epsilon)=\mathcal{N}(0,\sigma^{2}I_{n})$, which gives $\mathbb{E}[vv^{\top}]=\mathbb{E}\left[\frac{\epsilon}{||\epsilon||}\frac{\epsilon^{\top}}{||\epsilon||}\right]=\frac{1}{n}I_{n}$. ∎
Now, consider gradient descent in the direction of the gradient $\nabla f(\theta)$ as
$\theta_{t+1}=\theta_{t}+\alpha\nabla f(\theta_{t}).$ (32)
This update can be reformulated using directional derivatives.
Theorem A.4.
Let $\eta=\alpha n$ and $\epsilon^{(i)}\sim\mathcal{N}(0,\sigma^{2}I_{n})$. Gradient descent is equivalent to the update rule
$\theta_{t+1}=\theta_{t}+\alpha\sum_{i=1}^{K}\left[\frac{\epsilon^{(i)}}{||\epsilon^{(i)}||^{2}}\left[f\left(\theta+\epsilon^{(i)}\right)-f\left(\theta\right)\right]^{\top}\right]$ (33)
in the limit when $\sigma^{2}\to 0$ and $K\rightarrow\infty$.
Proof.
$\nabla f(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[v\nabla_{v}f(\theta)^{\top}\right]\,$
$=n\mathbb{E}_{p(\epsilon)}\left[v\left[\frac{f(\theta+h\epsilon)-f(\theta)}{h||\epsilon||}\,\right]^{\top}\right]\,$
$=n\mathbb{E}_{p(\epsilon)}\left[\frac{\epsilon}{||\epsilon||}\left[\frac{f(\theta+h\epsilon)-f(\theta)}{h||\epsilon||}\,\right]^{\top}\right]\,$
[…]
$\nabla f(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[\frac{\epsilon}{||\epsilon||^{2}}\left[f(\theta+\epsilon)-f(\theta)\,\right]^{\top}\right]\,$
We may approximate the expectation by sampling under some empirical distribution to obtain
$\nabla f(\theta)=n\sum_{i=1}^{K}\left[\frac{\epsilon^{(i)}}{||\epsilon^{(i)}||^{2}}\left[f\left(\theta+\epsilon^{(i)}\right)-f\left(\theta\right)\right]^{\top}\right]$
In practice, sufficiently small $\sigma^{2}$ and limited noise samples approximate the gradient. For $K=1$, analogous to single-sample updates in stochastic gradient descent:
$\theta_{t+1}=\theta_{t}+\eta\frac{\epsilon}{||\epsilon||^{2}}\delta f$ (34)
with $\delta f=\left[f(\theta+\epsilon)-f(\theta)\right]^{\top}$.
Appendix B Model hyperparameters and training details
All hyperparameters were carefully tuned per method and problem to ensure fair comparisons across methods. The number of episodes was chosen to illustrate the convergence of our method relative to baseline methods, with 8000, 20000, and 1000 episodes used for the Acrobot, Cartpole, and Reaching problems, respectively. In all our experiments, we set the smoothing factor for reward estimation to $\lambda=0.66$, striking a balance between recent values (for quick adaptability) and a longer history (for robustness against rapid reward fluctuations). Each neural network consisted of an input, hidden, and output layer. Input units corresponded to environment observation elements, and output units to possible actions. Specifically, Acrobot, Cartpole, and Reaching used 6, 4, and 32 input units and 3, 2, and 2 output units, respectively, with hidden layer sizes of 64, 64, and 128. Table 1 summarizes the learning rate $\eta$ and noise standard deviation $\sigma$ for each method and problem. Higher values of these parameters could lead to unstable training, while lower values may result in slower learning.
Table 1: Learning rate $\eta$ and noise standard deviation $\sigma$ for the different learning algorithms across problems. Dashes indicate that the parameter is not used.
BP Ours RMHL
[…]

### v2

@@
 Abstract
-Recent advances in reinforcement learning (RL) have led to significant improvements in task performance.
-However, training neural networks in an RL regime is typically achieved in combination with backpropagation, limiting their applicability in resource-constrained environments or when using non-differentiable neural networks.
-While noise-based alternatives like reward-modulated Hebbian learning (RMHL) have been proposed, their performance has remained limited, especially in scenarios with delayed rewards, which require retrospective credit assignment over time.
-Here, we derive a novel noise-based learning rule that addresses these challenges.
-Our approach combines directional derivative theory with Hebbian-like updates to enable efficient, gradient-free learning in RL.
-It features stochastic noisy neurons which can approximate gradients, and produces local synaptic updates modulated by a global reward signal.
-Drawing on concepts from neuroscience, our method uses reward prediction error as its optimization target to generate increasingly advantageous behavior, and incorporates an eligibility trace to facilitate temporal credit assignment in environments with delayed rewards.
-Its formulation relies on local information alone, making it compatible with implementations in neuromorphic hardware.
-Experimental validation shows that our approach significantly outperforms RMHL and is competitive with BP-based baselines, highlighting the promise of noise-based, biologically inspired learning for low-power and real-time applications.
+Biological neural systems efficiently learn from delayed rewards despite relying on noisy synaptic transmission and lacking centralized optimization mechanisms.
+In contrast, artificial neural networks trained with reinforcement learning typically rely on backpropagation (BP), which limits their use in resource-constrained systems or with non-differentiable components.
+While noise-based alternatives, like reward-modulated Hebbian learning (RMHL), provide a biologically grounded framework for credit assignment, they struggle with temporal delays and hierarchical processing —key challenges in real-world learning.
+In this work, we derive a novel noise-based learning rule to address these challenges.
+Drawing inspiration from biological neural circuits, our method uses reward prediction errors as its optimization target to generate increasingly advantageous behavior, and incorporates an eligibility trace to facilitate retrospective credit assignment.
+Its formulation relies on local information, aligning with biological constraints and enabling neuromorphic implementation.
+Experimental validation on reinforcement tasks (immediate and delayed rewards) shows our approach significantly outperforms RMHL and achieves performance comparable to BP, although with slower convergence due to its noise-driven updates.
+While tested on simple architectures, the results highlight the potential of noise-driven, brain-inspired learning for low-power adaptive systems, particularly in scenarios where energy efficiency and biological plausibility are a priority.
+These findings also offer mechanistic insights into how dopamine-like signals and synaptic stochasticity may jointly enable learning in biological networks, bridging computational models with neurobiological principles.
 1 Introduction
-Noise-based learning methods, which leverage stochastic fluctuations in neural activity [15, 20, 54, 52, 13] or synaptic weights [10, 57, 5, 8] to approximate gradients, offer a pathway toward simpler and more light-weight optimization techniques, that may reflect mechanisms within the brain itself [43, 29].
+The brain’s remarkable ability to learn and adapt to its environment is believed to arise from synaptic plasticity [4, 3, 37].
+These local modifications allow neural circuits to acquire complex sensorimotor skills despite inherent challenges such as noisy neural communication [12], temporally delayed feedback/reward [23], and no centralized error-propagation mechanism to drive weight updates [48].
+Understanding the mechanisms underlying this robust learning remains a central challenge in neuroscience and is increasingly inspiring novel approaches in artificial intelligence [11, 38].
+When it comes to learning in artificial neural networks, backpropagation (BP) underpins modern deep learning [45].
+However, its implementation in biological circuits is often seen as challenging [34].
+BP requires precise, non-local error gradients, which contrasts with the locality of synaptic plasticity that depends on pre- and post-synaptic activity and neuromodulatory factors —akin to three-factor learning [21, 29].
+Moreover, its deterministic nature is contrary to the inherent stochasticity of neural circuits [12, 51, 39], and its requirement for continuous, immediate error feedback contradicts biological learning, which often involves delayed feedback following a sequence of actions [33].
+While some works have investigated how variants of backpropagation might be implemented in cortical microcircuits —employing predictive coding, feedback alignment, or dendritic error signals to carry nonlocal gradients [60, 59, 31]— others have proposed different plasticity rules inspired directly from neurobiology [46, 25, 63].
+Among these approaches, noise‐based learning methods are especially promising.
+They leverage fluctuations in neural activity [19, 24, 61, 58, 16] or synaptic weights [13, 65, 7, 10], for estimating gradients.
+By injecting small random perturbations, these methods avoid the need for perfectly matched feedback pathways and rely only on information that biological synapses could plausibly access, aligning more closely with mechanisms observed in cortical neurons [50, 36].
 Given the prevalence of noise in biological neural systems [12, 51, 39], using noise as a mechanism for learning synaptic weights is an area of growing interest, with reward-modulated Hebbian learning (RMHL) [32, 40] as a promising candidate.
-RMHL, a form of three-factor Hebbian learning [17, 24], offers a potentially powerful mechanism for credit assignment without explicit backpropagation of exact gradients.
+RMHL, which is a form of three-factor Hebbian learning [21, 29], offers a potentially powerful mechanism for credit assignment without explicit backpropagation of errors.
 Despite theoretical advances, few noise-based methods have been adapted to real-world tasks.
@@
 This gap has limited the integration of noise-based approaches for effective credit assignment in temporally extended setups.
-In this work, we address this gap by proposing a novel noise-driven learning mechanism for reinforcement learning (RL), referred to as noise-based reward-modulated learning (NRL), which leverages noise to approximate gradients for synaptic updates derived from reward maximization.
-NRL integrates key concepts from neuroscience, such as eligibility traces [19], which link past neural states to future rewards, and reward prediction error (RPE) [41, 40, 34], which drives learning toward increasingly more advantageous behaviors.
-We employ directional derivatives for gradient approximation, which is realized through stochastic neurons that are compatible with the inherently noisy nature of brains and other physical systems.
-NRL estimates the directional derivative via two forward passes: a “noisy” pass with stochastic neurons and a “clean” pass without noise.
-In scenarios where a noiseless pass is infeasible, as in noisy analog hardware, multiple noisy passes can be averaged to approximate the clean pass, maintaining performance.
-We further validate our approach on RL benchmarks with both immediate- and delayed-reward problems.
-Finally, we discuss the broader implications for machine learning, neuromorphic computing, and neuroscience.
+To bridge these gaps, we propose noise-based reward-modulated learning (NRL), a biologically-inspired, gradient-free learning method compatible with delayed feedback/rewards.
+NRL builds on established neuroscientific pillars.
+It employs eligibility traces [23, 26], which tag synapses based on recent pre/post activity and noise perturbations, enabling retrospective credit assignment over behavioral timescales.
+It also embraces reward prediction error (RPE), that is, dopamine-like signals [48, 47, 42] that modulate these traces to reinforce actions that yield unexpected positive rewards, driving learning toward increasingly more advantageous behaviors.
+We derive our learning rule from first principles, employing directional derivatives to compute a local gradient estimate at each synapse, rather than relying on explicit gradients.
+These directional derivatives are implemented through stochastic neurons, aligning with the inherently noisy nature of neural circuits and other physical systems.
+NRL estimates the directional derivative with two forward passes: a “noisy” pass with stochastic neurons and a “clean” pass without noise.
+In scenarios where a noiseless pass is infeasible, multiple noisy passes can be averaged to approximate the clean pass, maintaining performance.
+We deliberately choose an abstract modelling framework so that our experiments isolate the learning rule’s behavior without introducing additional biological complexities.
+We employ rate-based neurons instead of spiking models, such that the dynamics reduce to smooth activation functions and continuous outputs.
+We also avoid distinguishing excitatory and inhibitory populations, allowing every unit to follow the same update equation.
+Finally, we treat potentiation and depression as two ends of a single weight-change continuum rather than separate processes, so that each synapse simply strengthens or weakens in proportion to its eligibility trace and reinforcement signal.
+In this way, we trade-off biological detail for more clear insight into capabilities and limitations of our proposed noise-based learning.
+We validate NRL on a suite of reinforcement-learning benchmarks spanning immediate-reward and delayed-reward tasks, demonstrating significantly superior learning efficiency when compared to RMHL (particularly when rewards were delayed by many steps), and competitive final performance compared to gradient-based baselines.
+Moreover, unlike RMHL, which breaks down in deeper architectures, NRL scales to multilayer networks, though converging more slowly due to its intrinsic stochasticity.
+Finally, we explore how these findings shed light on biological learning and point toward energy-efficient implementations on neuromorphic hardware, as well as novel, alternative, robust learning schemes for machine learning.
 2 Methods
-2.1 Agent-environment setup and learning signals
-We consider an agent interacting continuously with a dynamic environment, whose state at time $t$ is denoted by $S_{t}$.
-The agent observes $S_{t}$ and responds by selecting actions, $A_{t}$, which influence the environment.
-Each action or sequence of actions yields feedback in the form of positive or negative rewards $r_{t}$.
-This observe-act loop provides a feedback loop which can be leveraged to enable an agent to learn and adapt to the changing environment, aiming to maximize the rewards associated with its actions.
-The agent’s decision-making is guided by a policy $\pi$, representing the probability of choosing an action, $a$, in the state $S_{t}$, i.e., $\pi(a\mid S_{t})$.
-This policy reflects the agent’s understanding of optimal actions in each state.
-To determine an action, the agent computes a probability distribution over the set of possible actions, modeled as a categorical distribution, where $\sum_{a}\pi(a\mid S_{t})=1$, such that
-$P(A_{t}=a\mid S_{t})=\pi(a\mid S_{t})\,.$ (1)
-Rewards, $r_{t}$, from the environment, received after each action or sequence of actions, are used to compute the learning signal.
-Internally, the agent maintains a prediction of reward, denoted by $\bar{r}_{t}$.
-This prediction could be formed by an arbitrary computation, but for simplicity it is modeled as a running average over recent rewards
-$\bar{r}_{t}=\bar{r}_{t-1}+\lambda\left(r_{t}-\bar{r}_{t-1}\right)$ (2)
-where $\lambda$ is a smoothing factor. The mismatch between the predicted reward and the actual reward, denoted by
-$\delta r_{t}=r_{t}-\bar{r}_{t}$ (3)
-serves as the primary learning signal.
-It can be interpreted as a reward prediction error (RPE) in neuroscience contexts and can be thought of as a global dopaminergic signal that modulates learning in the brain [41, 40, 34].
-While conceptually similar to biological RPEs, this formulation is a simplification due to the straightforward running average used for reward prediction.
-Nonetheless, it effectively captures the core ideas of the proposed learning mechanism.
-In the Discussion, we explore how this prediction error could be refined.
+2.1 Neural system-environment interaction and reward signaling
+We model a neural system within a dynamic environment, grounded in the reinforcement learning framework [54].
+In this framework, the environment’s state at time $t$ is denoted by $s_{t}$.
+The neural system perceives $s_{t}$ and responds by initiating actions, $a_{t}$, which influence the environment.
+Each action or sequence of actions results in feedback in the form of positive or negative rewards $r_{t}$.
+This iterative process of perception, action, and outcome forms a feedback loop that enables the system to learn and adapt to the changing environment, ultimately seeking to maximize the rewards it receives.
+The system’s decision-making is guided by a policy $\pi$, representing the probability of choosing an action, $a_{t}$, in a given state $s_{t}$ at time $t$, i.e., $\pi(a_{t}\mid s_{t})$.
+This policy captures the system’s learned strategy for choosing actions that are likely to yield favorable outcomes.
+To determine an action, the system computes a probability distribution across the set of possible actions, modelled as a categorical distribution $P(a_{t}\mid s_{t})\equiv\pi(a_{t}\mid s_{t})$ with $\sum_{a\in A}\pi(a\mid s_{t})=1$, where $A$ denotes the set of all possible actions.
+The rewards $r_{t}$, obtained from the environment, serve as crucial learning signals.
+Internally, the system maintains a prediction of the expected reward, denoted by $\bar{r}_{t}$.
+While this prediction can be calculated through various mechanisms, we model it here for simplicity as a running average of recent rewards:
+$\bar{r}_{t+1}=\bar{r}_{t}+\lambda\left(r_{t+1}-\bar{r}_{t}\right)$ (1)
+where $\lambda$ is a smoothing factor that governs the influence of past rewards on the current prediction.
+The mismatch between the actual reward and the predicted reward, known as the reward prediction error, is then computed as
+$\delta_{t}=r_{t}-\bar{r}_{t}.$ (2)
+This RPE signal is a central element of our model and draws a strong parallel to the phasic activity of dopamine neurons in the brain [48, 47, 42].
+Dopamine neurons are known to exhibit increased activity for unexpected rewards and decreased activity for the unexpected absence of reward [17].
+Although our running average is a simplified representation of how reward expectations are formed in biological systems, it effectively captures the fundamental dynamics of the RPE and its role as a learning signal.
+We will explore potential refinements of this reward prediction mechanism in the Discussion section.
 2.2 Derivation of the learning rule
@@
 We begin by establishing a gradient-based learning rule rooted in an optimization target that aims to maximize unexpected rewards from the environment by increasing the reward prediction error.
-Maximizing the RPE rather than the direct reward enables the agent to adapt dynamically to changing environments, improve long-term performance by continuously seeking rewards that exceed expectations, and maintain robustness against variations in reward structures [48, 36].
-Building on this foundation, we transition to a gradient-free formulation by leveraging a directional derivative framework, where stochastic noise within the network approximates gradients.
+Using RPE, rather than the direct reward, enables our system to adapt dynamically to changing environments by tracking how rewards deviate from expectations, improving long-term performance by continuously seeking rewards that exceed expectations, and maintaining robustness against variations in reward structures [54, 43].
+Building on this foundation, we transition to a directional derivative framework, where stochastic noise within the network approximates gradients.
 This eliminates the need for explicit backpropagation and feedback phases, enabling learning through forward passes alone.
@@
 2.2.1 Gradient-based learning rule
-The agent is represented by a neural network which processes input observations and generates a probability distribution over possible actions.
-This neural network consists of multiple layers, each performing a linear transformation of its inputs, followed by a non-linear transformation as
-$x^{l}=f(h^{l})=f(W^{l}x^{l-1})$ (4)
-where $x^{l}$ represents the output of layer $l$, $h^{l}$ represents the layer pre-activation, $W^{l}$ is the weight matrix, and $f(\cdot)$ is a non-linear activation function.
-In this setup, the goal is to optimize the weights $W=\{W^{1},\ldots,W^{l}\}$ across the $L$ layers in the neural network to maximize the received rewards.
-Our derivation begins by introducing a generic variable $\theta$, which will later be mapped to the specific neural network parameters $W$.
-The primary learning objective is to maximize the expected RPE, $\delta r_{t}$, which aims to achieve higher-than-anticipated rewards.
-For now, let us consider learning to maximise RPE at a specific timepoint $t$.
-Since the reward prediction is a running average of recent rewards, the learning rule seeks to outperform previously received rewards.
-We express this objective in terms of the policy as
-$J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\delta r_{t}\right]\,.$ (5)
-To optimize $J(\theta)$, we can incrementally update the weights using the gradients with respect to $\theta$ as
-$\theta\leftarrow\theta+\eta\nabla_{\theta}J(\theta)$ (6)
+The system is modelled as a multi-layered network of interconnected units, analogous to populations of neurons, which processes input representing sensory observations and produces a probability distribution over possible actions.
+For simplicity, we refer to these populations as ‘layers,’ where each layer performs a transformation of its input in two stages: first, a linear transformation, followed by a non-linear transformation.
+This can be expressed mathematically as
+$x^{l}_{t}=f(h^{l}_{t})=f(W^{l}_{t}x^{l-1}_{t})$ (3)
+where $x^{l}_{t}$ represents the output activity of layer $l$, $h^{l}_{t}$ represents the layer pre-activation (the combined input to such layer), $W^{l}_{t}$ represents the matrix of synaptic weights connecting layer l1l-1 to layer $l$, at time $t$, and $f(\cdot)$ is a non-linear activation function, which introduces crucial non-linearities into the network’s computations
+In this setup, the goal is to adjust the synaptic weights, $\{W^{1},\ldots,W^{L}\}$, across the $L$ layers of the network to maximize the rewards obtained by the system.
+Our derivation begins by introducing a general parameter $\theta$, which will later be mapped to the specific neural network parameters $W$.
+The primary learning objective is to maximize the RPE, $\delta_{t}$, at any particular moment in time, $t$, effectively driving the system to seek out actions that lead to higher-than-predicted rewards.
+The reward prediction is modeled as a running average of recent rewards and, as such, the learning rule seeks to outperform previously received rewards.
+Note that for a fixed reward prediction, this is equivalent to maximizing the reward itself.
+We express this objective in terms of a parameterized policy at time $t$ as
+$J(\theta_{t})=\mathbb{E}_{\pi_{\theta_{t}}}\left[\delta_{t}\right]\,.$ (4)
+To optimize $J(\theta_{t})$, we can incrementally update the weights using the gradients with respect to $\theta_{t}$ as
+$\theta_{t+1}\leftarrow\theta_{t}+\eta\nabla J(\theta_{t})$ (5)
 where $\eta$ is the learning rate, controlling the size of each update step.
 Applying the policy gradient theorem [55], which utilizes the likelihood-ratio method, we express the gradient of the objective as
-$\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\nabla_{\theta}\log\pi_{\theta}\left(a\mid S_{t}\right)\delta r_{t}\right]\,.$ (7)
-For empirical estimation, in the case of a single sample, we approximate the gradient as
-$\nabla_{\theta}J(\theta)\approx\nabla_{\theta}\log\pi_{\theta}\left(a\mid S_{t}\right)\delta r_{t}\,.$ (8)
-With this approximation, we define the parameter updates as
-$\theta\leftarrow\theta+\eta\;\nabla_{\theta}\log\pi_{\theta}\left(a\mid S_{t}\right)\;\delta r_{t}.$ (9)
-Equation (9) resembles the REINFORCE update rule [54] but differs by using the reward prediction error (RPE), $\delta_{r}$, as the learning signal instead of cumulative rewards over full trajectories.
+$\nabla J(\theta_{t})=\mathbb{E}_{\pi_{\theta_{t}}}\left[\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)\delta_{t}\right]\,.$
+For empirical estimation, in the case of a single sample, we approximate the gradient as $\nabla J(\theta_{t})\approx\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)\delta_{t}$.
+Using this approximation, we define the parameter updates as
+$\theta_{t+1}\leftarrow\theta_{t}+\eta\;\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)\;\delta_{t}\,.$ (6)
+Equation (6) resembles the REINFORCE update rule [61] but differs by using the reward prediction error, $\delta_{t}$, as the learning signal instead of cumulative rewards over full trajectories.
 This RPE-based approach leverages immediate feedback from rewards as they are obtained rather than requiring a full trial completion to estimate the policy gradient.
-It shares conceptual similarities with actor-critic methods [1], where the policy is adjusted using a temporal difference error.
-However, we approximate future rewards with a running average of past rewards instead of a critic network, maintaining adaptive feedback benefits with a simpler implementation.
+It shares conceptual similarities with actor-critic methods in reinforcement learning [1], where the policy is adjusted using a temporal difference error.
+However, we approximate future rewards with a running average of past rewards instead of a critic network, a less powerful but simpler implementation which maintains adaptive feedback.
 2.2.2 Noise-based learning rule
 The learning rule derived in the previous section still relies on gradient descent.
-To avoid using backpropagation (BP) to compute the gradients, we propose a noise-based alternative that extends Equation (9).
-This approach leverages gradient approximation via directional derivatives, as detailed in Appendix A, enabling a theoretically rigorous derivation of noise-driven learning.
+To avoid using backpropagation to compute the gradients, we propose a noise-based alternative that extends Equation (6).
+This approach leverages gradient approximation via directional derivatives, enabling a theoretically rigorous derivation of noise-driven learning.
 A directional derivative quantifies the rate of change of a function in a specified direction.
-In neural networks, we implement this concept by introducing random noise into parameters.
+In our neural system, we implement this concept by introducing random noise into parameters.
 By comparing the network’s parameters with and without this noise, we obtain an estimate of the gradient direction.
-To formalize this, we express the gradient term in terms of directional derivatives using Theorem 31 in Appendix A as
-$\nabla p(\theta)$ $=\nabla_{\theta}\log\pi_{\theta}(a\mid S_{t})=n\mathbb{E}\left[v\nabla_{v}p(\theta)\right]$ (10)
-where $v=\ifrac{\epsilon}{||\epsilon||}$ is a normalized direction vector derived from noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$ with $n$ the number of parameters.
-We may expand the above via Theorem A.4, and add a difference term (parameters with noise versus without noise) to approximate the gradient by sampling under an empirical distribution
-$\nabla p(\theta)=n\sum_{i=1}^{K}\left[\frac{\epsilon^{(i)}}{||\epsilon^{(i)}||^{2}}\left(p\bigl{(}\tilde{\theta}^{(i)}\bigr{)}-p\bigl{(}\theta\bigr{)}\right)\right]\,.$ (11)
-where $K$ denotes the number of samples and $\tilde{\theta}^{(i)}=\theta+\epsilon^{(i)}$ the noise-perturbed parameters.
+To formalize this, we define $g(\theta_{t})=\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$ and express the gradient term in terms of directional derivatives using the theorems in Appendix A as
+$\nabla g(\theta_{t})$ $=\nabla\log\pi_{\theta_{t}}(a_{t}\mid s_{t})=n\mathbb{E}\left[\bar{\epsilon}_{t}\nabla_{\bar{\epsilon}_{t}}g(\theta_{t})\right]$ (7)
+[…]
+We may expand the above to approximate the gradient via a finite-difference and by sampling under an empirical distribution
+[…]
+where $K$ denotes the number of samples and $\tilde{\theta}_{t}^{(i)}=\theta_{t}+\epsilon^{(i)}_{t}$ the noise-perturbed parameters (see Appendix A).
 In practice, we consider $K=1$, analogous to single-sample updates in stochastic gradient descent.
-Here, $p\bigl{(}\tilde{\theta}\bigr{)}=\log\pi_{\tilde{\theta}}(a\mid S_{t})$ represents the log of the noise-perturbed output, while, $p\left(\theta\right)=\log\pi_{\theta}(a\mid S_{t})$ is the log of the noise-free output.
-Thus, $\rho=\log\pi_{\tilde{\theta}}(a\mid S_{t})-\log\pi_{\theta}(a\mid S_{t})$ captures the impact of the noise on the policy.
+Here, $g\bigl{(}\tilde{\theta}^{(i)}_{t}\bigr{)}=\log\pi_{\tilde{\theta}_{t}}(a_{t}\mid s_{t})$ and $g\left(\theta_{t}\right)=\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$ are the log of the noise-perturbed and noise-free output, respectively.
+Thus, we define $\rho_{t}=\log\pi_{\tilde{\theta}_{t}}(a_{t}\mid s_{t})-\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$, which captures the impact of the noise on the policy.
 Putting this together for the $K=1$ case, we obtain our noise-based learning rule
-$\theta\leftarrow\theta+\eta\delta_{r}\epsilon^{*}\rho$ (12)
-with $\epsilon^{*}=\ifrac{\epsilon}{||\epsilon||^{2}}$. For convenience, we absorbed the term $n$ into the learning rate $\eta$.
-We can extend this by introducing noise into the neurons instead of the weights of the network, allowing localized perturbations within network layers.
+$\theta_{t+1}\leftarrow\theta_{t}+\eta\delta_{t}\hat{\epsilon}_{t}\rho_{t}$
+[…]
+2.2.3 Learning with node level noise in neural network
+The noise-based learning rule derived so far is formulated for a general parameter $\theta_{t}$.
+Now, we apply this framework to the specific context of a neural network, where $\theta_{t}$ corresponds to the synaptic weights $W^{l}_{t}$ of layer $l$ at time $t$.
+Instead of directly perturbing the weights of the network $W^{l}_{t}$, we propose extending the learning rule by introducing noise directly into the neurons (nodes) of each layer.
+This strategy is advantageous because it can lead to reduced variance in gradient estimation, as perturbations occur in a lower-dimensional space (the neural activations) compared to the full weight space.
+This approach shares similarities with node perturbation (NP) [19, 24, 61, 58, 16], which has shown benefits in reducing gradient variance and offering more localized updates.
 Given a layer transformation as in Equation (3), adding noise at the neuron level is represented as
-$\tilde{x}^{l}=f\left(\tilde{h}^{l}+\xi^{l}\right)=f\left(W^{l}\tilde{x}^{l-1}+\xi^{l}\right)$ (13)
-where $\xi^{l}\sim\mathcal{N}(0,\beta\sigma^{2}I_{m^{l}})$ is the injected noise, $\beta$ is some arbitrarily small noise scale, and $m^{l}$ is the number of neurons in layer $l$.
-We use the notation $\tilde{x}^{l}$ and $\tilde{h}^{l}$ to denote perturbed inputs and pre-activations, respectively, which may also result from perturbations of previous network layers on which layer $l$ depends.
-To formulate the learning rule for specific layer parameters $W^{l}\in W$, we re-express the gradient in terms of the pre-activations instead of the parameters themselves.
+$\tilde{x}^{l}_{t}=f\left(\tilde{h}^{l}_{t}+\xi^{l}_{t}\right)=f\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (9)
+where $\xi^{l}_{t}\sim\mathcal{N}(0,\sigma^{2}I_{m^{l}})$ is the injected noise at time $t$, $\sigma^{2}$ is some arbitrarily small noise scale, and $m^{l}$ is the number of neurons in layer $l$.
+We use the notation $\tilde{x}^{l}_{t}$ and $\tilde{h}^{l}_{t}$ to denote perturbed inputs and pre-activations, respectively, which may also result from perturbations of previous network layers on which layer $l$ depends.
+To formulate the learning rule for specific layer parameters $W^{l}_{t}\in\{W^{1}_{t},\ldots,W^{L}_{t}\}$, we re-express the gradient in terms of the pre-activations instead of the parameters themselves.
+Here, the general parameter vector $\theta_{t}=\{W^{1}_{t},\ldots,W^{L}_{t}\}$ corresponds to the full set of all weight matrices for all layers at time $t$.
 To do so, we first rewrite the gradient term as
-$\nabla p\left(W^{l}\right)=\nabla_{W^{l}}\log\pi_{\theta}\left(a\mid S_{t}\right)$ (14)
-and apply the chain rule of differentiation to break down the gradient with respect to $W^{l}$ as
-$\nabla p\left(W^{l}\right)=\frac{\partial p}{\partial\tilde{h}^{l}}\frac{\partial\tilde{h}^{l}}{\partial W^{l}}=\nabla p\left(\tilde{h}^{l}\right)\tilde{x}^{l-1}\,.$ (15)
-Building on the theory developed above, we derive the gradient $\nabla p(\tilde{h})$, leading to the layer-specific weight update rule
-$W^{l}\leftarrow W^{l}+\eta\delta r_{t}\bar{\xi}^{l}{\rho}\tilde{x}^{l-1}$ (16)
-with ${\rho}=\log\pi_{\tilde{W}}(a\mid S_{t})-\log\pi_{W}(a\mid S_{t})$ and $\bar{\xi}^{l}=\ifrac{\xi^{l}}{||\xi^{l}||^{2}}$.
-Here, $\pi_{\tilde{W}}(a\mid S_{t})$ represents the network’s output when noise is injected into the neurons, while $\pi_{W}(a\mid S_{t})$ corresponds to the output of the noiseless network.
+$\nabla g\left(W^{l}_{t}\right)=\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)$ (10)
+and apply the chain rule to break down the gradient with respect to $W^{l}_{t}$ as
+$\nabla g\left(W^{l}_{t}\right)=\nabla g\left(h^{l}_{t}\right)\frac{\partial h^{l}_{t}}{\partial W^{l}_{t}}=\nabla g\left(h^{l}_{t}\right){{x}^{l-1\top}_{t}}.$ (11)
+Notably, we may once again carry out the conversion from this gradient estimation step to a set of directional derivatives such that
+[…]
+where $i$ indexes over a set of repeated samples of the noise term, $\xi$.
+Again, as above, by reducing this to the single sample ($n=1$) case, we can define a layer-specific weight update rule
+$W^{l}_{t+1}\leftarrow W^{l}_{t}+\eta\delta_{t}\bar{\xi}^{l}_{t}{\rho}_{t}{\tilde{x}^{l-1\top}_{t}}$ (13)
+[…]
 2.2.4 Learning from sparse rewards
-In our derivation above, synaptic updates occur at every time step, implying that learning signals and rewards are received continuously.
-However, rewards are usually provided after completing a sequence of actions or upon reaching specific milestones, leading to more sparsely spaced feedback.
-To handle this scenario, we modify our learning rule to accommodate sparse rewards arriving at arbitrary times.
-Starting from Equation (9), we redefine the synaptic weight update to include time explicitly
-$W^{l}_{t+1}=W^{l}_{t}+\eta\;\delta r_{t}\;\bar{\xi}^{l}_{t}p_{t}\tilde{x}^{l-1}_{t}$ (17)
-where $p_{t}=\log\pi_{\tilde{W}}(a\mid S_{t})-\log\pi_{W}(a\mid S_{t})$.
-Next, we assume sparse rewards arriving at arbitrary times $\tau$, and thus synaptic updates only occur at this moment, denoted by $W^{l}_{\tau}$.
-The learning rule now accumulates information since the last reward was received, updating weights by integrating this information over time
-$W^{l}_{\tau}\leftarrow W^{l}_{\tau_{0}}+\eta\;\delta r_{\tau}\;\sum_{t=\tau_{0}}^{\tau}\bar{\xi}^{l}_{t}{p}_{t}\tilde{x}^{l-1}_{t}$ (18)
-where $\tau_{0}$ represents the time when the previous reward was obtained.
-We further represent this accumulated sum using an eligibility trace, which captures the history of neural activity and noise contributions since the last reward, defined as
-$e_{\tau}=\sum_{t=\tau_{0}}^{\tau}\bar{\xi}^{l}_{t}{p}_{t}\tilde{x}^{l-1}_{t}\,.$ (19)
-such that (18) may be expressed as
-$W^{l}_{\tau}\leftarrow W^{l}_{\tau_{0}}+\eta\;\delta r_{\tau}\;e_{\tau}\,.$ (20)
-Thus, our final learning rule consists of two primary components: (i) an eligibility trace, defined in Equation (19), which accumulates local information over time at each time step, and (ii) a synaptic update, defined in Equation (20), triggered upon reward receipt, which modulates the eligibility trace to adjust the synaptic weights.
-Eligibility traces act as a mechanism to connect past actions with future rewards [19].
-They capture neural activity and other local variables over time, signaling potential synaptic changes.
+Our derivations so far assume that synaptic updates happen at every time step, implying a continuous stream of rewards and learning signals.
+In reality, however, rewards are often sparse, arriving only after a sequence of actions or upon reaching specific milestones.
+To handle this, we’ll modify our learning rule to account for rewards received at arbitrary times.
+Let’s denote these discrete reward times as $\tau_{m}\in\{\tau_{0},\tau_{1},...\tau_{M}\}$. At these moments, we compute the RPE, $\delta_{\tau_{m}}$, as
+$\delta_{\tau_{m}}=r_{\tau_{m}}-\bar{r}_{\tau_{m}}.$ (14)
+Here, $r_{\tau_{m}}$ is the reward received at time $\tau_{m}$, and $\bar{r}_{\tau_{m}}$ is the reward prediction, which is updated as the running average of recent rewards.
+$\bar{r}_{\tau_{m}}=\bar{r}_{\tau_{m-1}}+\lambda\left(r_{\tau_{m}}-\bar{r}_{\tau_{m-1}}\right)$ (15)
+While synaptic updates only occur at these specific reward times $\tau_{m}$, the learning rule continuously accumulates information between rewards.
+This information is integrated over time, allowing us to update weights based on what’s happened since the last reward.
+This leads to our modified learning rule
+$W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\;e_{\tau_{m}}$ (16)
+where
+$e_{\tau_{m}}=\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{l}_{t}{\rho}_{t}{\tilde{x}^{l-1\top}_{t}}$ (17)
+is an eligibility trace that acts as a mechanism to connect past actions with future rewards [23].
+Eligibility traces capture neural activity and other local variables over time, signaling potential synaptic changes.
 Upon receiving a reward, these traces are modulated by the reinforcement signal, resulting in synaptic updates.
@@
 The eligibility trace efficiently tracks neural information between rewards, facilitating the assignment of credit to past actions.
-2.3 Neural network setup
-At the core of the agent’s learning and decision-making process, we use a feedforward neural network with $L$ layers, where the transformation applied to each hidden layer is given by
-$\tilde{x}^{l}$ $=f\left(W^{l}\tilde{x}^{l-1}+\xi^{l}\right)$ (21)
+Thus, our final learning rule consists of two primary components: (i) an eligibility trace, defined in Equation (17), which accumulates local information over time at each time step, and (ii) a synaptic update, defined in Equation (16), triggered upon reward receipt, which modulates the eligibility trace to adjust the synaptic weights.
+This learning rule constitutes the core of our proposed noise-based reward-modulated Learning (NRL) and is used for all experiments presented in this paper.
+2.3 Neural network architecture
+The learning and decision-making capabilities of the system are modelled using a feedforward network with $L$ layers, representing interconnected populations of neurons.
+The transformation performed by each hidden layer at time $t$ can be expressed as
+$\tilde{x}^{l}_{t}$ $=f\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (18)
 and the readout in the output layer is given by
-$\tilde{y}$ $=s\left(W^{l}\tilde{x}^{l-1}+\xi^{l}\right)$ (22)
-where $a^{l}$ and $x^{l}$ denote the layer pre-activation and output (with $x_{0}$ as the network input), $y$ is the network output, $\xi^{l}\sim\mathcal{N}(0,\beta\sigma^{2}I_{m^{l}})$ is noise added to the neurons, with $m^{l}$ representing the number of units in layer $l$.
-This is referred to as a “noisy pass”.
-The function $f(\cdot)$ represents a non-linear activation function applied to each layer.
-In our case, we use the LeakyReLU activation function such that $f(x)=x$ if $x\geq 0$ and $f(x)=\alpha x$ with $0<\alpha\ll 1$, otherwise.
-The function $s(\cdot)$ represents a softmax transformation ${s}_{i}(x)=\ifrac{e^{x_{i}}}{\sum_{j=1}^{m^{l}}e^{x_{j}}}$, converting the network’s output into action probabilities.
-We also define a “clean pass,” where no noise is present in the network, as
-$x^{l}$ $=f\left(W^{l}x^{l-1}\right)$ (23)
-$y$ $=s\left(W^{l}x^{l-1}\right)$ (24)
+$\tilde{y}_{t}$ $=s\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (19)
+where $x^{l}_{t}$ represents the activity of units in layer $l$, $W^{l}_{t}$ is the matrix of synaptic weights connecting layer l1l-1 to layer $l$, and $\xi^{l}_{t}\sim\mathcal{N}(0,\sigma^{2}I_{m^{l}})$ represents random noise introduced into layer $l$, with $m^{l}$ being the number of units in such layer.
+This noisy propagation of activity is termed a “noisy pass.” The function $f(\cdot)$ is a non-linear activation function, which we implement as the LeakyReLU function such that $f(x)=x$ if $x\geq 0$ and $f(x)=\alpha x$ with $0<\alpha\ll 1$.
+The function $s(\cdot)$ represents a softmax transformation ${s}_{i}(x)=\ifrac{e^{x_{i}}}{\sum_{j=1}^{m^{l}}e^{x_{j}}}$, which transforms the network’s output into a probability distribution over possible actions.
+We also define a “clean pass,” representing the network’s activity in the absence of noise, by $x^{l}_{t}=f\left(W^{l}_{t}x^{l-1\top}_{t}\right)$ and $y_{t}=s\left(W^{l}_{t}x^{l-1\top}_{t}\right)$.
 2.4 Experimental validation
-We validate our approach across various environments and compare its performance with established baselines.
-Each environment presents an episodic problem, where the agent’s learning is guided solely by positive or negative rewards.
-We test in both instantaneous and delayed reward settings.
-In the delayed reward setting, a single reward is given after a sequence of actions, challenging the agent to assign credit retrospectively.
+We validate our approach across a range of simulated environments, comparing it against established baseline learning methods.
+Each environment presents a problem with discrete episodes, where the system’s learning is guided solely by positive or negative rewards.
+We investigate both scenarios with immediate reward and those involving delayed reward.
+In the delayed reward scenarios, a single reward is provided after a sequence of actions, challenging our system to assign credit retrospectively to the actions that contributed to the outcome.
 For the instantaneous reward setting, we use the Reaching problem [22], while the delayed reward setting uses the Cartpole [1] and Acrobot [53] problems.
 We use implementations given by the libraries OpenAI Gym [5] and NeuroGym [41] for the different environments.
-We compare NRL to two baselines: an exact-gradient version of NRL (similar to an actor-only variant of actor-critic which relies on a running average of past rewards as the prediction error), which serves as an ”optimal performance” benchmark and an RMHL approach.
-The first baseline uses BP for gradient computation, updating eligibility traces as
-$e_{\tau}=\sum_{t=\tau_{0}}^{\tau}\nabla_{W}\log\pi_{W}\left(a\mid S_{t}\right)\,$ (25)
+We compare NRL to two baselines: an exact-gradient version of NRL, which serves as an ”optimal performance” benchmark, and an RMHL approach.
+The first baseline, which we refer to as the exact-gradient method (BP baseline), is similar to an actor-only variant of actor-critic methods, relying on a running average of past rewards as the prediction error.
+For this baseline, we compute the exact policy gradients for the entire episode trajectory using backpropagation through time (BPTT) [57].
+Given sparse rewards, the weights at each layer are updated at the end of each episode according to
+$W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\left(\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})\right)\delta_{\tau_{m}}$ (20)
 The second baseline, motivated by the noise-based nature of NRL, is the RMHL rule from [32] with an explicit-noise approximation – the original version where noise is inferred from neural activities proved too unstable for the problems considered here.
 For delayed rewards, we adapt it similarly to [40], updating eligibility traces as
-$e_{\tau}=\sum_{t=\tau_{0}}^{\tau}\xi^{l}_{t}\tilde{x}^{l-1}_{t}\,.$ (26)
-In environments with high reward variability, like Cartpole and Acrobot, we stabilize synaptic updates by scaling the RPE by dividing by $r_{\tau}$.
+$e_{\tau_{m}}=\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{l}_{t}{\tilde{x}^{l-1\top}_{t}}\,.$ (21)
+Similar to NRL’s noise-based learning rule defined in Equation (16), the synaptic weights for RMHL are then updated by modulating this eligibility trace with the reward prediction error.
+In Table 1, we present the complete three learning rules (NRL, BPTT, and RMHL) in a single table for visualization and quick comparison.
+All updates are written in a per-layer form using $W^{l}$ and assume a delayed reward received at time $\tau_{m}$.
+Table 1: Comparison of the learning rules.
+Rule Update
+NRL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{\,l}_{t}\,\rho_{t}\,\tilde{x}^{\,l-1\top}_{t}$
+BPTT $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\left(\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})\right)\delta_{\tau_{m}}$
+RMHL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{\,l}_{t}\,\tilde{x}^{\,l-1\top}_{t}$
+In environments with high reward variability, like Cartpole and Acrobot, we stabilize synaptic updates by scaling the RPE by dividing by $r_{\tau_{m}}$.
 Hyperparameter values and training details are provided in Appendix B.
@@
 Training details are provided in Appendix B.
-In Section 3.2, we extend the comparison to deeper networks with multiple hidden layers.
-First, the Reaching problem [18], visualized in Figure 1C, is an instantaneous reward problem that requires an agent to reach and stay at a target on a 1D ring by moving left, right, or staying still.
-At each step, the agent observes both the target’s position and its own, and it is rewarded based on its proximity to the target over fixed-duration trials.
-Average across-trial and final performance (mean of the last 50 trials) are shown in Figures 1A and 1B, respectively.
+In Section titled “Scalability to multi-layer networks”, we extend the comparison to deeper networks with multiple hidden layers.
+First, the Reaching problem [22], visualized in Fig.
+1C, is an immediate reward problem that requires the system to reach and maintain a position at a target on a 1D ring by moving left, right, or remaining stationary.
+At each step, the system receives information about both the target’s position and its own, and a reward is provided based on proximity to the target over fixed-duration trials.
+Average performance across-trial and final performance (mean of the last 50 trials) are shown in Fig 1A and Fig 1B, respectively.
 Figure 1: Performance on benchmarks.
@@
 Right panels: Problem visualization.
-Second, the Acrobot problem [47], visualized in Figure 1F, involves delayed reward and requires controlling a two-link robotic arm to reach a target height.
-At each time step, the agent observes the angles and angular velocities of the two links and chooses one action: clockwise torque, counterclockwise torque, or no torque.
-Rewards are based on the speed of completion, with a maximum time.
-Average across-trial and final performance (mean of the last 50 trials) are shown in Figures 1D and 1E, respectively.
-The third and most challenging problem, the Cartpole problem [1], visualized in Figure 1I, is a delayed reward problem where an agent must balance a pole on a cart by moving left or right.
-At each time step, the agent observes the cart’s position and velocity, along with the pole’s angle and angular velocity, and acts accordingly.
-Performance is measured by the time the pole remains balanced, with a maximum time.
-Average across-trial and final performance (mean of the last 50 trials) are shown in Figures 1G and 1H, respectively.
-For all three tasks, we can observe that NRL reaches the performance level of BP whereas RMHL fails to do so.
-3.2 Scalability to deeper networks
-Figure 2: Performance on deeper networks – Acrobot.
+Second, the Acrobot problem [53], visualized in Fig 1F, involves delayed reward and requires controlling a two-link robotic arm to reach a target height.
+At each time step, the system receives information about the angles and angular velocities of the two links and chooses one action: clockwise torque, counterclockwise torque, or no torque.
+Rewards are given based on the speed of completion, with a maximum time allowed.
+Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig 1D and Fig 1E, respectively.
+The third and most challenging problem, the Cartpole problem [1], visualized in Fig 1I, is a delayed reward problem where the system must balance a pole on a cart by moving left or right.
+At each time step, the system receives information about the cart’s position and velocity, along with the pole’s angle and angular velocity, and responds accordingly.
+Performance is measured by the time the pole remains balanced, with a maximum time allowed.
+Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig 1G and Fig 1H, respectively.
+For all three tasks, NRL reaches the performance level of BP, whereas RMHL fails to do so.
+3.2 Scalability to multi-layer networks
+Here, we demonstrate that NRL can effectively assign credit in neural networks with multiple hidden layers; a challenging scenario where most biologically plausible algorithms struggle.
+We use both the Acrobot and the Cartpole tasks for this purpose, as they present a delayed reward problem, making them a more realistic test for credit assignment in reinforcement learning.
+In this comparison, we include the same baselines, BP and RMHL.
+Fig 2 displays the results for the Acrobot task and Fig 3 displays the results for the Cartpole tasks.
+The left panels of these two figures show performance across trials, averaged over 5 runs.
+In contrast, the right panels display the final performance (mean of the last 50 trials), also averaged over 5 runs.
+The results for neural networks with two hidden layers are shown in Fig 2A and Fig 2B, and 3A and B, and those for networks with three hidden layers are presented in Fig 2C and Fig 2D, and Fig 3C and Fig 3D.
+Figure 2: Performance on deeper networks for the Acrobot.
 A, B: 2-hidden layer networks.
@@
 Right panels: Final performance (mean of the last 50 trials), averaged over 5 runs.
-Figure 3: Performance on deeper networks – Cartpole.
+Figure 3: Performance on deeper networks for the Cartpole.
 A, B: 2-hidden layer networks.
@@
 Right panels: Final performance (mean of the last 50 trials), averaged over 5 runs.
-Here, we demonstrate that NRL can effectively assign credit in neural networks with multiple hidden layers – a challenging scenario where most biologically plausible algorithms struggle.
-We use both the Acrobot and the Cartpole tasks for this purpose, as they present a delayed reward problem, making them a more realistic test for credit assignment in reinforcement learning.
-In this comparison, we include the same baselines, BP and RMHL.
-Figure 2 displays the results for the Acrobot task and Figure 3 displays the results for the Cartpole tasks.
-The left panels of these two figures show performance across trials, averaged over 5 runs.
-In contrast, the right panels display the final performance (mean of the last 50 trials), also averaged over 5 runs.
-The results for neural networks with two hidden layers are shown in Figures 2A and B, and 3A and B, and those for networks with three hidden layers are presented in Figures 2C and D, and 3C and D.
 Our results indicate that NRL successfully learns to solve the tasks, achieving performance comparable to BP.
 In contrast, RMHL struggles with credit assignment in deeper networks.
-However, NRL requires more trials to converge as the network depth increases—an expected outcome due to the stochastic nature of the updates [20].
+However, NRL requires more trials to converge as the network depth increases – an expected outcome due to the stochastic nature of the updates [24].
 A similar trend is observed with BP, though to a lesser extent, as its gradient-based updates inherently provide more directed adjustments.
 3.3 Using only noisy passes
-In NRL, ${p}_{t}$ is calculated as the difference in the network’s output between the clean and noisy passes.
+In NRL, $\rho_{t}$ is calculated as the difference in the network’s output between the clean and noisy passes.
 However, a completely noiseless network is not feasible in more biophysically realistic scenarios.
-Instead, the clean network’s output can be approximated by averaging the outputs from multiple noisy passes, as the injected noise averages to zero over time $\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\xi_{i}^{l}=0$.
+Instead, the clean network’s output can be approximated by averaging the outputs from multiple noisy passes, as the injected noise averages to zero $\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\xi_{t,i}^{l}=0$.
 This assumes that the network dynamics are faster than the environment dynamics, allowing the network to perform multiple forward passes before the environment changes.
-Figure 4: Learning using only noisy passes – Acrobot.
+To evaluate the accuracy of this approximation, we employ a single hidden-layer network and the Acrobot problem.
+The difference between the clean pass output and the averaged noisy passes output was calculated for each timestep and averaged over 500 timesteps, as shown in Fig 4A.
+Furthermore, Fig 4B illustrates the performance on the Acrobot problem using only noisy forward passes, starting with a minimum of 2 noisy passes.
+We also extended this evaluation to 10 noisy passes to explore the impact of increasing the number of passes, showing slightly faster initial convergence.
+Figure 4: Learning using only noisy passes for the Acrobot.
 A: Clean pass approximation error.
@@
 Performance across trials, averaged over 5 runs, with mean, minimum, and maximum values displayed.
-To evaluate the accuracy of this approximation, we employ a single hidden-layer network and the Acrobot problem.
-The difference between the clean pass output and the averaged noisy passes output was calculated for each timestep and averaged over 500 timesteps, as shown in Figure 4A.
-Furthermore, Figure 4B illustrates the performance on the Acrobot problem using only noisy forward passes, starting with the minimum of 2 noisy passes.
-We also extended this evaluation to 10 noisy passes to explore the impact of increasing the number of passes, showing slightly faster initial convergence.
 4 Discussion
-In this work, we introduced NRL as a novel noise-driven learning mechanism for neural networks that bridges RL principles with Hebbian-like updates, providing a gradient-free alternative to traditional approaches.
-By leveraging directional derivatives and stochastic noisy neurons to approximate gradients [44, 53, 52], our method produces local synaptic updates that are modulated by a global reinforcement signal.
-This noise-based rule offers an efficient approach for neural adaptation, particularly suited to real-time and resource-limited environments.
+In this work, we propose NRL, a noise-driven synaptic plasticity rule that bridges principles of reinforcement learning and optimization with biologically-inspired Hebbian updates.
+By leveraging stochastic neural activity, our method produces local synaptic updates modulated by a global reinforcement signal.
+It offers an efficient approach for neural adaptation, particularly suited to neuromorphic systems.
+We employ a top-down methodology, starting by deriving our learning rule from the mathematical principles of gradient-based optimization.
+We then approximate the gradients via directional derivatives and incorporate bottom-up constraints inspired by biological systems.
+Specifically, the gradient-approximation process is implemented through stochastic neural activity (aligning with the inherent noise in the brain [12, 51, 39]), and reinforcement learning mechanisms are realized through reward prediction errors (emulating dopaminergic reward signals [48, 47]), eligibility traces, and local synaptic plasticity.
+These elements enable learning under uncertainty [18].
+This approach results in a learning rule that is both theoretically grounded and potentially realizable in neural circuits.
+Our results demonstrate the learning effectiveness of NRL, solving both immediate and delayed feedback/reward problems in simulated environments, achieving performance comparable to baselines optimized employing backpropagation.
+Additionally, it shows a significant improvement over RMHL [32, 40], a modern, well-known approach that employs noise as the core of its synaptic plasticity.
+Our scalability results further demonstrate that NRL effectively assigns credit even in multilayer networks, a challenge for many biologically plausible algorithms, indicating its potential for scaling.
+This is a critical difference with RMHL, which struggles to assign credit in multilayer networks evidenced by its minimal performance improvements across trials.
+Nevertheless, as the depth of the network increases, NRL shows slower convergence due to its noise-driven updates.
 NRL shares similarities with node perturbation methods [19, 24, 61, 58, 16], which inject noise into neurons to adjust weights based on state and performance fluctuations.
-However, unlike node perturbation –which requires constant feedback or reward for learning– NRL is well-suited for delayed-reward setups, a common scenario in real-world applications.
-By leveraging eligibility traces [19], NRL can effectively assign credit to past neural states even when actions and outcomes are separated in time.
-Furthermore, NRL shares conceptual similarities with forward gradient methods [2], which also estimate gradients using directional derivatives, stochastic perturbations, and forward-mode differentiation.
-Both frameworks leverage noise-driven updates, focusing on solving real-world tasks.
-However, while forward gradient methods rely solely on random perturbations to approximate gradients, NRL refines updates by incorporating reinforcement learning principles, using RPEs to directly modulate synaptic weights.
-Recent advances in forward gradient techniques, such as employing local auxiliary networks to produce informed gradient guesses [16], have demonstrated substantial improvements in alignment with exact gradients, narrowing the performance gap with backpropagation.
-Similarly, these variance-reduction strategies could inspire further refinements in our approach, enhancing the quality of noise-based updates and improving scalability to more complex architectures.
-Another increasingly popular approach for gradient estimation during optimization is zeroth-order (ZO) optimization [28, 6].
-Like NRL, common ZO optimization relies on perturbation-driven strategies, utilizing differences in function values resulting from input perturbations to approximate gradients.
-However, as with forward gradient methods, NRL distinguishes itself by incorporating elements, such as RPEs and eligibility traces, to learn in delayed reward scenarios.
-Unlike forward gradient methods, which rely on directional derivatives and access to the model’s structure, ZO methods estimate gradients purely from function evaluations, without requiring any knowledge of the network’s internal structure or differentiability.
-This makes ZO methods particularly well-suited for black-box optimization tasks, a flexibility that our approach also shares.
-The results demonstrate the effectiveness of NRL, significantly outperforming the comparable RMHL approach [26, 32] in both instantaneous and delayed reward tasks.
-While RMHL methods aim to align neural dynamics with observed brain dynamics rather than optimizing learning, our learning rule is rigorously derived to perform gradient approximation.
-As a result, our learning rule introduces two key innovations.
-First, normalizing the noise, $\bar{\xi}^{l}_{t}=\ifrac{\xi^{l}_{t}}{||\xi^{l}_{t}||^{2}}$, allows for smaller noise magnitudes that minimally impact agent behavior while still contributing effectively to the learning process.
-Second, an additional term, ${p}_{t}=\left[\log\pi_{\tilde{\theta}}(a|S_{t})-\log\pi_{\theta}(a|S_{t})\right]$, captures the influence of noise on the network output, thereby enhancing credit assignment accuracy.
-Our results further demonstrate that NRL effectively assigns credit even in deep networks with multiple hidden layers, a challenge for many biologically plausible algorithms.
-As network depth increases, NRL’s performance advantage over RMHL becomes more pronounced.
-In fact, RMHL struggles with credit assignment in networks with more than one hidden layer, as evidenced by its minimal performance improvements during training.
-Nevertheless, NRL does require more trials to converge in deeper networks, which is an expected consequence of its stochastic update mechanism [20].
-Compared to traditional BP-based methods, NRL approximates gradients rather than relying on exact gradients [27, 51, 39, 38], offering a significant advantage in computationally constrained scenarios—such or when using non-differentiable networks like spiking neural networks [55, 50].
-The local-update structure of NRL provides compatibility with neuromorphic hardware [42], where efficient, low-power computations are essential.
+However, unlike node perturbation – which requires constant feedback or reward for learning – NRL is well-suited for delayed reward setups.
+By leveraging eligibility traces [23, 26], it can effectively assign credit to past neural states even when actions and outcomes are separated in time.
+Gradient approximation, a cornerstone of our learning rule’s derivation, is also crucial in mathematical optimization and machine learning, enabling learning when exact gradients are computationally expensive or unavailable.
+Techniques such as forward gradient methods [2] and zeroth-order (ZO) optimization [35, 8] have emerged, utilizing stochastic perturbations and function value comparisons to estimate gradients.
+Recent advances, such as employing local auxiliary networks for informed gradient guesses [20], have significantly improved the alignment of these approximations with exact gradients, reducing the performance gap with backpropagation.
+These variance-reduction strategies could similarly enhance our approach by improving the quality of the updates and scalability to more complex architectures.
+Furthermore, compared to traditional gradient-based methods like backpropagation, gradient approximation methods offer a key advantage in computationally constrained scenarios, or when using non-differentiable networks, such as spiking neural networks [62, 56].
+Additionally, the local-update structure of NRL provides compatibility with neuromorphic substrates [49], where efficient, low-power computations are essential.
 Neuromorphic systems can also inherently simulate noisy neural activity, enabling energy-efficient learning in applications where real-time adaptation is key.
-The inherent noise in NRL is also advantageous for RL, as it naturally facilitates exploration-exploitation balance [22, 25].
+The inherent noise in NRL is also advantageous for reinforcement learning setups, as it naturally facilitates exploration-exploitation balance [27, 30].
 Injected noise perturbs neural states, promoting exploratory deviations from expected values and thus introducing variations in the agent’s policy.
-Our work further draws inspiration from neuroscience and underscores the importance of biologically plausible approaches [56, 30] for advancing learning algorithms.
-The reliance on RPEs as learning signals emulates the dopaminergic reward signals observed in the brain [41, 40], crucial for synaptic plasticity.
-The noise-based rule, coupled with eligibility traces, also aligns with biological processes that enable learning under uncertainty [14] and reflect the noisy nature of the brain [9, 45, 31].
-Notice that the definition of RPEs used in this work is a simplification, relying on a straightforward running average for reward prediction.
-While this is sufficient to capture the core ideas of the proposed learning mechanism, more sophisticated approaches could further enhance reward estimation.
-One such extension is Temporal Difference (TD) learning [46, 48], which provides a more principled method for predicting expected future rewards and aligns more closely with biological dopaminergic activity [35].
-A natural implementation would be an actor-critic architecture, where a critic network estimates the value function (e.g., expected future rewards), refining the reward prediction error (RPE).
-One limitation of NRL is its dependence on multiple forward passes per synaptic update, which, though parallelizable, reduces biological plausibility.
-Future work could explore efficient ways to estimate noise impacts without explicitly needing separate clean and noisy passes.
-Additionally, while BP may offer faster convergence in unrestricted computational settings, recent research, including this study and others [7, 11], show promise in bridging this gap for specific networks and tasks.
-Our findings suggest that noise-based learning strategies could drive advances in machine learning by enabling efficient, local, and energy-conscious algorithms, particularly suited for neuromorphic hardware and real-time applications.
-This research also offers valuable insights into how learning might occur in the brain, enriching our understanding of both artificial and biological intelligence.
+Our work has limitations that suggest future research directions.
+First, its dependence on multiple forward passes per synaptic update reduces biological plausibility, even though these passes could be performed in parallel.
+Future work should explore more efficient ways to estimate noise impacts, potentially eliminating the need for multiple passes.
+Second, our modeling approach deliberately involves a high degree of abstraction: we use rate-based neurons instead of spiking models, do not distinguish between excitatory and inhibitory populations, and treat synaptic potentiation and depression as a single continuum.
+While this provides clear insight into the capabilities and limitations of our approach, more biologically detailed models could offer further insights.
+Third, our reward prediction error calculation is a simplification.
+While sufficient for demonstrating the learning mechanism, the simplicity of using a running average may limit its effectiveness in scenarios where navigating through transient low-reward regions is necessary to achieve higher long-term rewards.
+In such cases, the model may underestimate the value of actions that lead to delayed benefits, potentially resulting in suboptimal updates.
+More sophisticated approaches, such as temporal difference learning [52, 54], or learned prediction through a critic network, could provide a richer and more temporally sensitive estimate of expected rewards.
+These strategies, which align more closely with biological dopaminergic signaling and can be implemented within actor-critic architectures, represent promising directions for improving learning performance in future work.
+Finally, while backpropagation may offer faster convergence in unconstrained settings, recent studies [9, 14], including ours, indicate that this gap can be narrowed for specific networks and tasks.
+Our findings highlight the potential of unifying optimization methods and neurobiological principles.
+By drawing inspiration from both domains, we have developed a learning rule that not only offers a biologically plausible alternative to backpropagation but also provides insights into how learning, particularly synaptic plasticity, might occur in biological neural systems.
+Conversely, the constraints of biological realism can guide the development of more efficient and robust artificial learning algorithms, particularly suited for neuromorphic hardware and real-time applications.
+This synergistic approach promises to advance our understanding of intelligence in both biological and artificial systems [64, 38], driving progress in neuroscience and machine learning alike.
 Acknowledgments
@@
 Appendix A Gradient approximation using directional derivatives
-Consider a function $f(\theta)$, where $\theta=(W_{1},\ldots,W_{L})$ are its parameters. The gradient of this function is defined as follows:
+Consider a function $g(\theta)$, where $\theta=(W_{1},\ldots,W_{L})$ are its parameters. The gradient of this function is defined as follows:
 Definition A.1.
-The gradient $\nabla f$ is a vector indicating the direction of the steepest ascent of the function $f$, with components as partial derivatives of $f(\theta)$:
-$\nabla f=\left[\frac{\partial f}{\partial\theta}\right]^{\top}=\left[\frac{\partial f}{\partial W_{1}},\ldots,\frac{\partial f}{\partial W_{L}}\right]^{\top}\,.$ (27)
-While $\nabla f$ captures the rate of change of $f$ in the steepest direction, a directional derivative gives the rate of change in a specified direction.
-For a unit vector $v=\ifrac{\epsilon}{||\epsilon||}$, normalized via the Euclidean norm, we define the directional derivative as:
+The gradient $\nabla g$ is a vector indicating the direction of the steepest ascent of the function $g$, with components as partial derivatives of $g(\theta)$:
+$\nabla g=\left[\frac{\partial g}{\partial\theta}\right]^{\top}=\left[\frac{\partial g}{\partial W_{1}},\ldots,\frac{\partial g}{\partial W_{L}}\right]^{\top}\,.$ (22)
+[…]
 Definition A.2.
-The directional derivative of $f(\theta)$ along a unit vector $v=(v_{1},\ldots,v_{n})$ is defined by the limit
-$\nabla_{v}{f}(\theta)=\lim_{h\to 0}{\frac{f(\theta+hv)-f(\theta)}{h}}\,,$ (28)
+The directional derivative of $g(\theta)$ along a unit vector $v=(v_{1},\ldots,v_{n})$ is defined by the limit
+$\nabla_{v}{g}(\theta)=\lim_{h\to 0}{\frac{g(\theta+hv)-g(\theta)}{h}}\,$, (23)
 where $h$ is a small step size.
-For a sufficiently small $h$ we approximate the directional derivative as
-$\nabla_{{v}}{f}(\theta)=\frac{f(\theta+h\epsilon)-f(\theta)}{h||\epsilon||}\,.$ (29)
-This directional derivative is essentially a projection of $\nabla f$ in the direction $v$, following the relation:
-$\nabla_{v}{f}(\theta)=v\cdot\nabla{f}(\theta).$ (30)
+For a sufficiently small $h$ we numerically measure the directional derivative as
+[…]
+This directional derivative can also be measured as a projection of $\nabla g$ in the direction $v$, following the relation:
+$\nabla_{v}{g}(\theta)=v\cdot\nabla{g}(\theta).$ (25)
 We can now formally demonstrate how gradients can be approximated using directional derivatives.
 Theorem A.3.
-Let $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$, drawn from some probability distribution $p(\epsilon)$, and $n$ the number of dimensions in $\theta$.
+Let $\epsilon\sim\mathcal{N}(0,\sigma^{2}I_{n})$ (probability distribution $p(\epsilon)$) where $n$ is the number of dimensions in $\theta$.
 Exact gradients can be written in terms of directional derivatives using expectations
-$\nabla f(\theta)=n\mathbb{E}_{p(\epsilon)}\left[v\nabla_{v}f(\theta)^{\top}\right]\,.$ (31)
+$\nabla g(\theta)=n\mathbb{E}_{p(\epsilon)}\left[v^{\top}\nabla_{v}g(\theta)\right]\,.$ (26)
 Proof.
-$\nabla f(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[v\nabla_{v}f(\theta)^{\top}\right]\,$
-$=n\mathbb{E}_{p(\epsilon)}\left[vv\nabla f(\theta)^{\top}\right]\,$
-$=n\mathbb{E}_{p(\epsilon)}\left[vv\right]\nabla f(\theta)^{\top}\,$
-$=n\mathbb{E}_{p(\epsilon)}\left[\frac{\epsilon}{||\epsilon||}\frac{\epsilon^{\top}}{||\epsilon||}\right]\nabla f(\theta)\,$
-$=n\frac{1}{n}I_{n}\nabla f(\theta)=\nabla f(\theta)\,$
-as we assume $p(\epsilon)=\mathcal{N}(0,\sigma^{2}I_{n})$, which gives $\mathbb{E}[vv^{\top}]=\mathbb{E}\left[\frac{\epsilon}{||\epsilon||}\frac{\epsilon^{\top}}{||\epsilon||}\right]=\frac{1}{n}I_{n}$.
-∎
-Now, consider gradient descent in the direction of the gradient $\nabla f(\theta)$ as
-$\theta_{t+1}=\theta_{t}+\alpha\nabla f(\theta_{t}).$ (32)
+$\nabla g(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[v^{\top}\nabla_{v}g(\theta)\right]\,$
+$=n\mathbb{E}_{p(\epsilon)}\left[vv^{\top}\nabla g(\theta)\right]\,$
+$=n\mathbb{E}_{p(\epsilon)}\left[vv^{\top}\right]\nabla g(\theta)\,$
+[…]
+$=n\frac{1}{n}I_{n}\nabla g(\theta)=\nabla g(\theta)\,$
+[…]
+Now, consider gradient descent in the direction of the gradient $\nabla g(\theta)$ as
+$\theta_{t+1}=\theta_{t}+\alpha\nabla g(\theta_{t}).$ (27)
 This update can be reformulated using directional derivatives.
 Theorem A.4.
-Let $\eta=\alpha n$ and $\epsilon^{(i)}\sim\mathcal{N}(0,\sigma^{2}I_{n})$. Gradient descent is equivalent to the update rule
-$\theta_{t+1}=\theta_{t}+\alpha\sum_{i=1}^{K}\left[\frac{\epsilon^{(i)}}{||\epsilon^{(i)}||^{2}}\left[f\left(\theta+\epsilon^{(i)}\right)-f\left(\theta\right)\right]^{\top}\right]$ (33)
-in the limit when $\sigma^{2}\to 0$ and $K\rightarrow\infty$.
+Let $\eta=\frac{\alpha n}{hK}$ and $\epsilon^{(i)}\sim\mathcal{N}(0,\sigma^{2}I_{n})$. Gradient descent is equivalent to the update rule
+[…]
+in the limit when $h\to 0$ and $K\rightarrow\infty$.
 Proof.
-$\nabla f(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[v\nabla_{v}f(\theta)^{\top}\right]\,$
-$=n\mathbb{E}_{p(\epsilon)}\left[v\left[\frac{f(\theta+h\epsilon)-f(\theta)}{h||\epsilon||}\,\right]^{\top}\right]\,$
-$=n\mathbb{E}_{p(\epsilon)}\left[\frac{\epsilon}{||\epsilon||}\left[\frac{f(\theta+h\epsilon)-f(\theta)}{h||\epsilon||}\,\right]^{\top}\right]\,$
+$\nabla g(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[v\nabla_{v}g(\theta)^{\top}\right]\,$
 […]
-$\nabla f(\theta)$ $=n\mathbb{E}_{p(\epsilon)}\left[\frac{\epsilon}{||\epsilon||^{2}}\left[f(\theta+\epsilon)-f(\theta)\,\right]^{\top}\right]\,$
-We may approximate the expectation by sampling under some empirical distribution to obtain
-$\nabla f(\theta)=n\sum_{i=1}^{K}\left[\frac{\epsilon^{(i)}}{||\epsilon^{(i)}||^{2}}\left[f\left(\theta+\epsilon^{(i)}\right)-f\left(\theta\right)\right]^{\top}\right]$
-In practice, sufficiently small $\sigma^{2}$ and limited noise samples approximate the gradient. For $K=1$, analogous to single-sample updates in stochastic gradient descent:
-$\theta_{t+1}=\theta_{t}+\eta\frac{\epsilon}{||\epsilon||^{2}}\delta f$ (34)
-with $\delta f=\left[f(\theta+\epsilon)-f(\theta)\right]^{\top}$.
+[…]
+[…]
+Now, substituting the expectation with a sampling under some empirical distribution,
+[…]
+Finally, defining gradient descent on parameters $\theta$ as $\theta_{t+1}=\theta_{t}+\alpha\nabla g(\theta)$,
+$\theta_{t+1}$ $=\theta_{t}+\alpha\nabla g(\theta)$
+[…]
+[…]
+In practice, small perturbations, controlled by $h$, and limited numbers of noise samples are sufficient to approximate the gradient.
+For $K=1$, analogous to single-sample updates in stochastic gradient descent:
+[…]
+with $\rho=\left[g(\theta+h\epsilon)-g(\theta)\right]^{\top}$.
 Appendix B Model hyperparameters and training details
@@
 BP Ours RMHL
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Acrobot $\eta$ 5e-3 5e-2 5e-2
+$\sigma$ - 1e-3 1e-3
+Cartpole $\eta$ 5e-3 5e-2 1e-2
+$\sigma$ - 1e-3 1e-1
+Reaching $\eta$ 1e-2 1e-2 1e-1
+$\sigma$ - 1e-3 1e-1

### v3

@@
 Abstract
-Biological neural systems efficiently learn from delayed rewards despite relying on noisy synaptic transmission and lacking centralized optimization mechanisms.
-In contrast, artificial neural networks trained with reinforcement learning typically rely on backpropagation (BP), which limits their use in resource-constrained systems or with non-differentiable components.
-While noise-based alternatives, like reward-modulated Hebbian learning (RMHL), provide a biologically grounded framework for credit assignment, they struggle with temporal delays and hierarchical processing —key challenges in real-world learning.
-In this work, we derive a novel noise-based learning rule to address these challenges.
-Drawing inspiration from biological neural circuits, our method uses reward prediction errors as its optimization target to generate increasingly advantageous behavior, and incorporates an eligibility trace to facilitate retrospective credit assignment.
-Its formulation relies on local information, aligning with biological constraints and enabling neuromorphic implementation.
-Experimental validation on reinforcement tasks (immediate and delayed rewards) shows our approach significantly outperforms RMHL and achieves performance comparable to BP, although with slower convergence due to its noise-driven updates.
-While tested on simple architectures, the results highlight the potential of noise-driven, brain-inspired learning for low-power adaptive systems, particularly in scenarios where energy efficiency and biological plausibility are a priority.
-These findings also offer mechanistic insights into how dopamine-like signals and synaptic stochasticity may jointly enable learning in biological networks, bridging computational models with neurobiological principles.
+The pursuit of energy-efficient and adaptive artificial intelligence (AI) has positioned neuromorphic computing as a promising alternative to conventional computing.
+However, achieving learning on these platforms requires techniques that prioritize local information while enabling effective credit assignment.
+Here, we propose noise-based reward-modulated learning (NRL), a novel synaptic plasticity rule that mathematically unifies reinforcement learning and gradient-based optimization with biologically-inspired local updates.
+NRL addresses the computational bottleneck of exact gradients by approximating them through stochastic neural activity, transforming the inherent noise of biological and neuromorphic substrates into a functional resource.
+Drawing inspiration from biological learning, our method uses reward prediction errors as its optimization target to generate increasingly advantageous behavior, and eligibility traces to facilitate retrospective credit assignment.
+Experimental validation on reinforcement tasks, featuring immediate and delayed rewards, shows that NRL achieves performance comparable to baselines optimized using backpropagation, although with slower convergence, while showing significantly superior performance and scalability in multi-layer networks compared to reward-modulated Hebbian learning (RMHL), the most prominent similar approach.
+While tested on simple architectures, the results highlight the potential of noise-driven, brain-inspired learning for low-power adaptive systems, particularly in computing substrates with locality constraints.
+NRL offers a theoretically grounded paradigm well-suited for the event-driven characteristics of next-generation neuromorphic AI.
 1 Introduction
-The brain’s remarkable ability to learn and adapt to its environment is believed to arise from synaptic plasticity [4, 3, 37].
-These local modifications allow neural circuits to acquire complex sensorimotor skills despite inherent challenges such as noisy neural communication [12], temporally delayed feedback/reward [23], and no centralized error-propagation mechanism to drive weight updates [48].
-Understanding the mechanisms underlying this robust learning remains a central challenge in neuroscience and is increasingly inspiring novel approaches in artificial intelligence [11, 38].
-When it comes to learning in artificial neural networks, backpropagation (BP) underpins modern deep learning [45].
-However, its implementation in biological circuits is often seen as challenging [34].
-BP requires precise, non-local error gradients, which contrasts with the locality of synaptic plasticity that depends on pre- and post-synaptic activity and neuromodulatory factors —akin to three-factor learning [21, 29].
-Moreover, its deterministic nature is contrary to the inherent stochasticity of neural circuits [12, 51, 39], and its requirement for continuous, immediate error feedback contradicts biological learning, which often involves delayed feedback following a sequence of actions [33].
-While some works have investigated how variants of backpropagation might be implemented in cortical microcircuits —employing predictive coding, feedback alignment, or dendritic error signals to carry nonlocal gradients [60, 59, 31]— others have proposed different plasticity rules inspired directly from neurobiology [46, 25, 63].
-Among these approaches, noise‐based learning methods are especially promising.
-They leverage fluctuations in neural activity [19, 24, 61, 58, 16] or synaptic weights [13, 65, 7, 10], for estimating gradients.
-By injecting small random perturbations, these methods avoid the need for perfectly matched feedback pathways and rely only on information that biological synapses could plausibly access, aligning more closely with mechanisms observed in cortical neurons [50, 36].
-Given the prevalence of noise in biological neural systems [12, 51, 39], using noise as a mechanism for learning synaptic weights is an area of growing interest, with reward-modulated Hebbian learning (RMHL) [32, 40] as a promising candidate.
+Modern artificial intelligence (AI) models are growing exponentially, leading to proportional increases in their computational demands and energy consumption [44].
+This trend, combined with the growing requirement for widespread AI deployment in local edge devices [52, 60], creates a massive demand for systems that are both fast and energy-efficient.
+Meeting this demand requires a fundamental shift in computing architecture to overcome the energy and latency bottlenecks of conventional Von Neumann systems [59].
+Neuromorphic computing [50, 45, 29] offers a promising path forward by mimicking the low energy, event-driven, and parallel processing capabilities of the brain [50, 35].
+Although neuromorphic hardware is highly efficient for inference, the ability to perform effective on-chip learning and adaptation remains the primary challenge, lacking suitable on-chip learning algorithms.
+Backpropagation (BP) remains the workhorse algorithm for training artificial neural networks, computing precise global gradients for weight updates [46].
+However, its operational requirements are fundamentally incompatible with the constraints of neuromorphic hardware [10].
+BP requires precise global error propagation, which translates into high-bandwidth global data movement and large memory requirements to store activations, the main source of energy consumption in deep learning systems [36].
+This contrasts sharply with the hardware needs for local synaptic updates that rely only on information available at the synaptic weight.
+In addition, BP’s iterative forward and backward passes are incompatible with the asynchronous event-driven regime in which most neuromorphic devices operate [41], and its deterministic nature cannot tolerate inherit noise in neuromorphic circuits.
+In response, local plasticity approaches and biologically-inspired learning rules have been proposed to bypass the constraints of BP [64, 63, 32, 47, 24, 67].
+A particularly promising family of methods is noise-based learning, which leverages stochastic perturbations (either in weights [12, 69, 5, 9] or activations [18, 23, 65, 62, 15]) to approximate the true direction of the gradient.
+By injecting small random perturbations, these methods avoid the need for perfectly matched feedback pathways and can leverage inherit noise [51, 35].
+Given the prevalence of noise on neuromorphic and biological substrates [11, 53, 37], using noise as a mechanism to learn synaptic weights is an area of growing interest, with reward-modulated Hebbian learning (RMHL) [33, 38] as a promising candidate.
 RMHL, which is a form of three-factor Hebbian learning [20, 30], offers a potentially powerful mechanism for credit assignment without explicit backpropagation of errors.
@@
 This gap has limited the integration of noise-based approaches for effective credit assignment in temporally extended setups.
-To bridge these gaps, we propose noise-based reward-modulated learning (NRL), a biologically-inspired, gradient-free learning method compatible with delayed feedback/rewards.
-NRL builds on established neuroscientific pillars.
-It employs eligibility traces [23, 26], which tag synapses based on recent pre/post activity and noise perturbations, enabling retrospective credit assignment over behavioral timescales.
-It also embraces reward prediction error (RPE), that is, dopamine-like signals [48, 47, 42] that modulate these traces to reinforce actions that yield unexpected positive rewards, driving learning toward increasingly more advantageous behaviors.
-We derive our learning rule from first principles, employing directional derivatives to compute a local gradient estimate at each synapse, rather than relying on explicit gradients.
-These directional derivatives are implemented through stochastic neurons, aligning with the inherently noisy nature of neural circuits and other physical systems.
-NRL estimates the directional derivative with two forward passes: a “noisy” pass with stochastic neurons and a “clean” pass without noise.
-In scenarios where a noiseless pass is infeasible, multiple noisy passes can be averaged to approximate the clean pass, maintaining performance.
-We deliberately choose an abstract modelling framework so that our experiments isolate the learning rule’s behavior without introducing additional biological complexities.
-We employ rate-based neurons instead of spiking models, such that the dynamics reduce to smooth activation functions and continuous outputs.
-We also avoid distinguishing excitatory and inhibitory populations, allowing every unit to follow the same update equation.
-Finally, we treat potentiation and depression as two ends of a single weight-change continuum rather than separate processes, so that each synapse simply strengthens or weakens in proportion to its eligibility trace and reinforcement signal.
-In this way, we trade-off biological detail for more clear insight into capabilities and limitations of our proposed noise-based learning.
-We validate NRL on a suite of reinforcement-learning benchmarks spanning immediate-reward and delayed-reward tasks, demonstrating significantly superior learning efficiency when compared to RMHL (particularly when rewards were delayed by many steps), and competitive final performance compared to gradient-based baselines.
+To bridge these gaps, we propose noise-based reward-modulated learning (NRL), a biologically-inspired, gradient-free learning method, which is compatible with delayed feedback/rewards.
+We derive our learning rule from first principles, employing directional derivatives to compute a local gradient estimate at each synapse.
+These directional derivatives are implemented through stochastic neurons, aligning with the noisy nature of neuromorphic physical systems, and are estimated with two forward passes: a “noisy” pass with stochastic neurons and a “clean” pass without noise.
+In scenarios where a noiseless pass is infeasible, multiple noisy passes can be averaged to approximate the clean pass, maintaining performance as shown in our experiments.
+NRL draws from established neuroscientific concepts, by employing eligibility traces [22, 25], which tag synapses based on recent pre/post activity and noise perturbations, allowing retrospective credit assignment over behavioral timescales.
+It also incorporates reward prediction errors (RPEs), a dopamine-like signal [49, 48, 40] that modulates eligibility traces to reinforce actions that yield unexpected positive rewards, driving learning toward increasingly more advantageous behaviors.
+We deliberately utilize a rate-based neural model, instead of a spiking model, to focus the analysis on the core computational capabilities and limitations of the learning rule itself, making our findings immediately applicable to neuromorphic substrates.
+We validate NRL on a suite of reinforcement-learning benchmarks that span immediate-reward and delayed-reward tasks, demonstrating significantly superior learning efficiency compared to RMHL (particularly when rewards were delayed by many steps) and competitive final performance compared to the gradient-based baseline.
 Moreover, unlike RMHL, which breaks down in deeper architectures, NRL scales to multilayer networks, though converging more slowly due to its intrinsic stochasticity.
-Finally, we explore how these findings shed light on biological learning and point toward energy-efficient implementations on neuromorphic hardware, as well as novel, alternative, robust learning schemes for machine learning.
+Our findings argue that NRL represents a promising and scalable step toward developing low-energy and fast adaptive AI systems.
+By transforming the stochasticity of neuromorphic hardware from a challenge into a computational resource, NRL provides a crucial new algorithmic path for enabling complex, continuous learning at the edge, as well as novel, alternative, robust learning schemes for machine learning.
 2 Methods
@@
 We model a neural system within a dynamic environment, grounded in the reinforcement learning framework [56].
-In this framework, the environment’s state at time $t$ is denoted by $s_{t}$.
-The neural system perceives $s_{t}$ and responds by initiating actions, $a_{t}$, which influence the environment.
+In this framework, the state of the environment at time $t$ is denoted by $s_{t}$.
+The neural system perceives $s_{t}$ and responds by initiating actions $a_{t}$, which influence the environment.
 Each action or sequence of actions results in feedback in the form of positive or negative rewards $r_{t}$.
 This iterative process of perception, action, and outcome forms a feedback loop that enables the system to learn and adapt to the changing environment, ultimately seeking to maximize the rewards it receives.
-The system’s decision-making is guided by a policy $\pi$, representing the probability of choosing an action, $a_{t}$, in a given state $s_{t}$ at time $t$, i.e., $\pi(a_{t}\mid s_{t})$.
+The decision-making of the system is guided by a policy $\pi$, which represents the probability of choosing an action, $a_{t}$, in a given state $s_{t}$ at time $t$, i.e., $\pi(a_{t}\mid s_{t})$.
 This policy captures the system’s learned strategy for choosing actions that are likely to yield favorable outcomes.
-To determine an action, the system computes a probability distribution across the set of possible actions, modelled as a categorical distribution $P(a_{t}\mid s_{t})\equiv\pi(a_{t}\mid s_{t})$ with $\sum_{a\in A}\pi(a\mid s_{t})=1$, where $A$ denotes the set of all possible actions.
+To determine an action, the system computes a probability distribution across the set of possible actions, modeled as a categorical distribution $P(a_{t}\mid s_{t})\equiv\pi(a_{t}\mid s_{t})$ with $\sum_{a\in A}\pi(a\mid s_{t})=1$, where $A$ denotes the set of all possible actions.
 The rewards $r_{t}$, obtained from the environment, serve as crucial learning signals.
 Internally, the system maintains a prediction of the expected reward, denoted by $\bar{r}_{t}$.
-While this prediction can be calculated through various mechanisms, we model it here for simplicity as a running average of recent rewards:
+Although this prediction can be calculated through various mechanisms, we model it here for simplicity as a running average of recent rewards as
 $\bar{r}_{t+1}=\bar{r}_{t}+\lambda\left(r_{t+1}-\bar{r}_{t}\right)$ (1)
@@
 $\delta_{t}=r_{t}-\bar{r}_{t}.$ (2)
-This RPE signal is a central element of our model and draws a strong parallel to the phasic activity of dopamine neurons in the brain [48, 47, 42].
-Dopamine neurons are known to exhibit increased activity for unexpected rewards and decreased activity for the unexpected absence of reward [17].
-Although our running average is a simplified representation of how reward expectations are formed in biological systems, it effectively captures the fundamental dynamics of the RPE and its role as a learning signal.
+This RPE signal, which draws a strong parallel to the phasic activity of dopamine neurons in the brain [49, 48, 40, 16], functions as a low-bandwidth global feedback signal that is inherently compatible with the communication constraints of neuromorphic hardware.
+Although our running average is a simplified representation of how reward expectations are formed, it effectively captures the fundamental dynamics of the RPE and its role as an adaptive learning signal.
 We will explore potential refinements of this reward prediction mechanism in the Discussion section.
@@
 We begin by establishing a gradient-based learning rule rooted in an optimization target that aims to maximize unexpected rewards from the environment by increasing the reward prediction error.
-Using RPE, rather than the direct reward, enables our system to adapt dynamically to changing environments by tracking how rewards deviate from expectations, improving long-term performance by continuously seeking rewards that exceed expectations, and maintaining robustness against variations in reward structures [54, 43].
-Building on this foundation, we transition to a directional derivative framework, where stochastic noise within the network approximates gradients.
-This eliminates the need for explicit backpropagation and feedback phases, enabling learning through forward passes alone.
+Using RPE, rather than direct reward, enables our system to adapt dynamically to changing environments by tracking how rewards deviate from expectations, improving long-term performance by continuously seeking rewards that exceed expectations, and maintaining robustness against variations in reward structures [56, 42].
+Building on this foundation, we transition to a directional derivative framework, where intrinsic noise within the network is leveraged to approximate gradients.
+This critical step eliminates the need for backpropagation and feedback phases, enabling learning through forward passes alone.
 Finally, we extend our noise-based approach to handle scenarios with delayed rewards, a hallmark of real-world problems.
-This extension enables NRL to adapt to environments where feedback is only obtained after a series of actions.
+This extension enables NRL to adapt to environments where feedback is obtained only after a series of actions.
 2.2.1 Gradient-based learning rule
-The system is modelled as a multi-layered network of interconnected units, analogous to populations of neurons, which processes input representing sensory observations and produces a probability distribution over possible actions.
+The system is modeled as a multi-layered network of interconnected units, analogous to populations of neurons, which processes input representing sensory observations and produces a probability distribution over possible actions.
 For simplicity, we refer to these populations as ‘layers,’ where each layer performs a transformation of its input in two stages: first, a linear transformation, followed by a non-linear transformation.
@@
 Now, we apply this framework to the specific context of a neural network, where $\theta_{t}$ corresponds to the synaptic weights $W^{l}_{t}$ of layer $l$ at time $t$.
-Instead of directly perturbing the weights of the network $W^{l}_{t}$, we propose extending the learning rule by introducing noise directly into the neurons (nodes) of each layer.
-This strategy is advantageous because it can lead to reduced variance in gradient estimation, as perturbations occur in a lower-dimensional space (the neural activations) compared to the full weight space.
+Instead of directly perturbing the weights of the network $W^{l}_{t}$, we propose introducing noise directly into the neurons (nodes) of each layer.
+This strategy is advantageous because it leads to reduced variance in gradient estimation, as perturbations occur in a lower-dimensional space (the neural activations) compared to the full weight space.
+In addition, it naturally reduces the computational cost and communication overhead associated with high-dimensional perturbations.
 This approach shares similarities with node perturbation (NP) [18, 23, 65, 62, 15], which has shown benefits in reducing gradient variance and offering more localized updates.
 Given a layer transformation as in Equation (3), adding noise at the neuron level is represented as
-$\tilde{x}^{l}_{t}=f\left(\tilde{h}^{l}_{t}+\xi^{l}_{t}\right)=f\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (9)
+$\tilde{x}^{l}_{t}=f\left(\tilde{h}^{l}_{t}+\xi^{l}_{t}\right)=f\left(W^{l}_{t}(\tilde{x}^{l-1}_{t})^{\top}+\xi^{l}_{t}\right)$ (9)
 where $\xi^{l}_{t}\sim\mathcal{N}(0,\sigma^{2}I_{m^{l}})$ is the injected noise at time $t$, $\sigma^{2}$ is some arbitrarily small noise scale, and $m^{l}$ is the number of neurons in layer $l$.
@@
 and apply the chain rule to break down the gradient with respect to $W^{l}_{t}$ as
-$\nabla g\left(W^{l}_{t}\right)=\nabla g\left(h^{l}_{t}\right)\frac{\partial h^{l}_{t}}{\partial W^{l}_{t}}=\nabla g\left(h^{l}_{t}\right){{x}^{l-1\top}_{t}}.$ (11)
+$\nabla g\left(W^{l}_{t}\right)=\nabla g\left(h^{l}_{t}\right)\frac{\partial h^{l}_{t}}{\partial W^{l}_{t}}=\nabla g\left(h^{l}_{t}\right)(x^{l-1}_{t})^{\top}.$ (11)
 Notably, we may once again carry out the conversion from this gradient estimation step to a set of directional derivatives such that
@@
 Again, as above, by reducing this to the single sample ($n=1$) case, we can define a layer-specific weight update rule
-$W^{l}_{t+1}\leftarrow W^{l}_{t}+\eta\delta_{t}\bar{\xi}^{l}_{t}{\rho}_{t}{\tilde{x}^{l-1\top}_{t}}$ (13)
+$W^{l}_{t+1}\leftarrow W^{l}_{t}+\eta\delta_{t}\bar{\xi}^{l}_{t}{\rho}_{t}(\tilde{x}^{l-1}_{t})^{\top}$ (13)
 […]
@@
 To handle this, we’ll modify our learning rule to account for rewards received at arbitrary times.
-Let’s denote these discrete reward times as $\tau_{m}\in\{\tau_{0},\tau_{1},...\tau_{M}\}$. At these moments, we compute the RPE, $\delta_{\tau_{m}}$, as
+Let us denote these discrete reward times as $\tau_{m}\in\{\tau_{0},\tau_{1},...\tau_{M}\}$. At these moments, we compute the RPE as
 $\delta_{\tau_{m}}=r_{\tau_{m}}-\bar{r}_{\tau_{m}}.$ (14)
-Here, $r_{\tau_{m}}$ is the reward received at time $\tau_{m}$, and $\bar{r}_{\tau_{m}}$ is the reward prediction, which is updated as the running average of recent rewards.
-$\bar{r}_{\tau_{m}}=\bar{r}_{\tau_{m-1}}+\lambda\left(r_{\tau_{m}}-\bar{r}_{\tau_{m-1}}\right)$ (15)
+Here, $r_{\tau_{m}}$ is the reward received at time $\tau_{m}$, and $\bar{r}_{\tau_{m}}$ is the reward prediction, which is updated as the running average of recent rewards:
+$\bar{r}_{\tau_{m}}=\bar{r}_{\tau_{m-1}}+\lambda\left(r_{\tau_{m}}-\bar{r}_{\tau_{m-1}}\right).$ (15)
 While synaptic updates only occur at these specific reward times $\tau_{m}$, the learning rule continuously accumulates information between rewards.
@@
 where
-$e_{\tau_{m}}=\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{l}_{t}{\rho}_{t}{\tilde{x}^{l-1\top}_{t}}$ (17)
+$e_{\tau_{m}}=\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{l}_{t}{\rho}_{t}(\tilde{x}^{l-1}_{t})^{\top}$ (17)
 is an eligibility trace that acts as a mechanism to connect past actions with future rewards [22].
@@
 The transformation performed by each hidden layer at time $t$ can be expressed as
-$\tilde{x}^{l}_{t}$ $=f\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (18)
+$\tilde{x}^{l}_{t}$ $=f\left(W^{l}_{t}(\tilde{x}^{l-1}_{t})^{\top}+\xi^{l}_{t}\right)$ (18)
 and the readout in the output layer is given by
-$\tilde{y}_{t}$ $=s\left(W^{l}_{t}\tilde{x}^{l-1\top}_{t}+\xi^{l}_{t}\right)$ (19)
+$\tilde{y}_{t}$ $=s\left(W^{l}_{t}(\tilde{x}^{l-1}_{t})^{\top}+\xi^{l}_{t}\right)$ (19)
 where $x^{l}_{t}$ represents the activity of units in layer $l$, $W^{l}_{t}$ is the matrix of synaptic weights connecting layer l1l-1 to layer $l$, and $\xi^{l}_{t}\sim\mathcal{N}(0,\sigma^{2}I_{m^{l}})$ represents random noise introduced into layer $l$, with $m^{l}$ being the number of units in such layer.
@@
 The function $s(\cdot)$ represents a softmax transformation ${s}_{i}(x)=\ifrac{e^{x_{i}}}{\sum_{j=1}^{m^{l}}e^{x_{j}}}$, which transforms the network’s output into a probability distribution over possible actions.
-We also define a “clean pass,” representing the network’s activity in the absence of noise, by $x^{l}_{t}=f\left(W^{l}_{t}x^{l-1\top}_{t}\right)$ and $y_{t}=s\left(W^{l}_{t}x^{l-1\top}_{t}\right)$.
+We also define a “clean pass,” representing the network’s activity in the absence of noise, by $x^{l}_{t}=f\left(W^{l}_{t}(x^{l-1}_{t})^{\top}\right)$ and $y_{t}=s\left(W^{l}_{t}(x^{l-1}_{t})^{\top}\right)$.
 2.4 Experimental validation
@@
 Given sparse rewards, the weights at each layer are updated at the end of each episode according to
-$W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\left(\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})\right)\delta_{\tau_{m}}$ (20)
+$W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\left(\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})\right)\delta_{\tau_{m}}.$ (20)
 The second baseline, motivated by the noise-based nature of NRL, is the RMHL rule from [33] with an explicit-noise approximation – the original version where noise is inferred from neural activities proved too unstable for the problems considered here.
 For delayed rewards, we adapt it similarly to [38], updating eligibility traces as
-$e_{\tau_{m}}=\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{l}_{t}{\tilde{x}^{l-1\top}_{t}}\,.$ (21)
+$e_{\tau_{m}}=\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{l}_{t}(\tilde{x}^{l-1}_{t})^{\top}\,.$ (21)
 Similar to NRL’s noise-based learning rule defined in Equation (16), the synaptic weights for RMHL are then updated by modulating this eligibility trace with the reward prediction error.
@@
 All updates are written in a per-layer form using $W^{l}$ and assume a delayed reward received at time $\tau_{m}$.
-Table 1: Comparison of the learning rules.
+Table 1: Learning rules. The terms in the BPTT update are rearranged for easier comparison.
 Rule Update
-NRL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{\,l}_{t}\,\rho_{t}\,\tilde{x}^{\,l-1\top}_{t}$
-BPTT $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\left(\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})\right)\delta_{\tau_{m}}$
-RMHL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{\,l}_{t}\,\tilde{x}^{\,l-1\top}_{t}$
+NRL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{\,l}_{t}\,\rho_{t}\,(\tilde{x}^{l-1}_{t})^{\top}$
+BPTT $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\delta_{\tau_{m}}\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})$
+RMHL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{\,l}_{t}\,(\tilde{x}^{l-1}_{t})^{\top}$
 In environments with high reward variability, like Cartpole and Acrobot, we stabilize synaptic updates by scaling the RPE by dividing by $r_{\tau_{m}}$.
@@
 3 Results
-3.1 Solving instantaneous and delayed reward problems
-To ensure comparability with RMHL methods, which typically utilize single-hidden-layer networks, we first conduct experiments with a one-hidden-layer neural network for BP, NRL, and RMHL.
-Training details are provided in Appendix B.
-In Section titled “Scalability to multi-layer networks”, we extend the comparison to deeper networks with multiple hidden layers.
-First, the Reaching problem [22], visualized in Fig.
-1C, is an immediate reward problem that requires the system to reach and maintain a position at a target on a 1D ring by moving left, right, or remaining stationary.
-At each step, the system receives information about both the target’s position and its own, and a reward is provided based on proximity to the target over fixed-duration trials.
-Average performance across-trial and final performance (mean of the last 50 trials) are shown in Fig 1A and Fig 1B, respectively.
+3.1 Performance on control tasks with instantaneous and delayed rewards
 Figure 1: Performance on benchmarks.
@@
 Right panels: Problem visualization.
-Second, the Acrobot problem [53], visualized in Fig 1F, involves delayed reward and requires controlling a two-link robotic arm to reach a target height.
+To ensure comparability with RMHL methods, which typically utilize single-hidden-layer networks, we first conduct experiments with a one-hidden-layer neural network for BP, NRL, and RMHL.
+Training details are provided in Appendix B.
+In Section titled “Scalability to deeper architectures”, we extend the comparison to deeper networks with multiple hidden layers.
+First, the Reaching problem [21], visualized in Fig.
+1C, is an immediate reward problem that requires the system to reach and maintain a position at a target on a 1D ring by moving left, right, or remaining stationary.
+At each step, the system receives information about both the target’s position and its own, and a reward is provided based on proximity to the target over fixed-duration trials.
+Average performance across-trial and final performance (mean of the last 50 trials) are shown in Fig.
+1A and Fig.
+1B, respectively.
+Second, the Acrobot problem [55], visualized in Fig.
+1F, involves delayed reward and requires controlling a two-link robotic arm to reach a target height.
 At each time step, the system receives information about the angles and angular velocities of the two links and chooses one action: clockwise torque, counterclockwise torque, or no torque.
 Rewards are given based on the speed of completion, with a maximum time allowed.
-Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig 1D and Fig 1E, respectively.
-The third and most challenging problem, the Cartpole problem [1], visualized in Fig 1I, is a delayed reward problem where the system must balance a pole on a cart by moving left or right.
+Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig.
+1D and Fig.
+1E, respectively.
+The third and most challenging problem, the Cartpole problem [1], visualized in Fig.
+1I, is a delayed reward problem where the system must balance a pole on a cart by moving left or right.
 At each time step, the system receives information about the cart’s position and velocity, along with the pole’s angle and angular velocity, and responds accordingly.
 Performance is measured by the time the pole remains balanced, with a maximum time allowed.
-Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig 1G and Fig 1H, respectively.
-For all three tasks, NRL reaches the performance level of BP, whereas RMHL fails to do so.
-3.2 Scalability to multi-layer networks
+Average performance across trials and final performance (mean of the last 50 trials) are shown in Fig.
+1G and Fig.
+1H, respectively.
+For all three tasks, NRL achieves a final performance comparable to the gradient-based baseline, BP, demonstrating a vast difference compared to RMHL.
+3.2 Scalability to deeper architectures
 Here, we demonstrate that NRL can effectively assign credit in neural networks with multiple hidden layers; a challenging scenario where most biologically plausible algorithms struggle.
@@
 In this comparison, we include the same baselines, BP and RMHL.
-Fig 2 displays the results for the Acrobot task and Fig 3 displays the results for the Cartpole tasks.
+Figure 2 displays the results for the Acrobot task and Fig.
+3 displays the results for the Cartpole tasks for neural networks consisting of two and three hidden layers.
 The left panels of these two figures show performance across trials, averaged over 5 runs.
 In contrast, the right panels display the final performance (mean of the last 50 trials), also averaged over 5 runs.
-The results for neural networks with two hidden layers are shown in Fig 2A and Fig 2B, and 3A and B, and those for networks with three hidden layers are presented in Fig 2C and Fig 2D, and Fig 3C and Fig 3D.
 Figure 2: Performance on deeper networks for the Acrobot.
@@
 In contrast, RMHL struggles with credit assignment in deeper networks.
-However, NRL requires more trials to converge as the network depth increases – an expected outcome due to the stochastic nature of the updates [24].
+However, NRL requires more trials to converge as the network depth increases, which is an expected outcome due to the stochastic nature of the updates [23].
 A similar trend is observed with BP, though to a lesser extent, as its gradient-based updates inherently provide more directed adjustments.
-3.3 Using only noisy passes
+3.3 Learning using only noisy passes
 In NRL, $\rho_{t}$ is calculated as the difference in the network’s output between the clean and noisy passes.
-However, a completely noiseless network is not feasible in more biophysically realistic scenarios.
-Instead, the clean network’s output can be approximated by averaging the outputs from multiple noisy passes, as the injected noise averages to zero $\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\xi_{t,i}^{l}=0$.
+However, generating a perfectly noiseless pass may be hardware-infeasible or energy-prohibitive in neuromorphic circuits.
+Instead, we show that the clean network’s output can be approximated by averaging the outputs from multiple noisy passes, as the injected noise averages to zero $\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\xi_{t,i}^{l}=0$.
 This assumes that the network dynamics are faster than the environment dynamics, allowing the network to perform multiple forward passes before the environment changes.
 To evaluate the accuracy of this approximation, we employ a single hidden-layer network and the Acrobot problem.
-The difference between the clean pass output and the averaged noisy passes output was calculated for each timestep and averaged over 500 timesteps, as shown in Fig 4A.
-Furthermore, Fig 4B illustrates the performance on the Acrobot problem using only noisy forward passes, starting with a minimum of 2 noisy passes.
+The difference between the clean pass output and the averaged noisy passes output was calculated for each timestep and averaged over 500 timesteps, as shown in Fig.
+4A.
+Furthermore, Fig.
+4B illustrates the performance on the Acrobot problem using only noisy forward passes, starting with a minimum of 2 noisy passes.
 We also extended this evaluation to 10 noisy passes to explore the impact of increasing the number of passes, showing slightly faster initial convergence.
@@
 4 Discussion
-In this work, we propose NRL, a noise-driven synaptic plasticity rule that bridges principles of reinforcement learning and optimization with biologically-inspired Hebbian updates.
+In this work, we propose NRL, noise-based reward-modulated learning, a novel synaptic plasticity rule that bridges the principles of reinforcement learning and optimization with biologically inspired Hebbian updates.
 By leveraging stochastic neural activity, our method produces local synaptic updates modulated by a global reinforcement signal.
 It offers an efficient approach for neural adaptation, particularly suited to neuromorphic systems.
-We employ a top-down methodology, starting by deriving our learning rule from the mathematical principles of gradient-based optimization.
-We then approximate the gradients via directional derivatives and incorporate bottom-up constraints inspired by biological systems.
-Specifically, the gradient-approximation process is implemented through stochastic neural activity (aligning with the inherent noise in the brain [12, 51, 39]), and reinforcement learning mechanisms are realized through reward prediction errors (emulating dopaminergic reward signals [48, 47]), eligibility traces, and local synaptic plasticity.
-These elements enable learning under uncertainty [18].
-This approach results in a learning rule that is both theoretically grounded and potentially realizable in neural circuits.
+We employ a top-down methodology, deriving our learning rule from the mathematical principles of gradient-based optimization.
+We approximate the gradients through directional derivatives and incorporate bottom-up constraints inspired by biological systems.
+Specifically, the gradient-approximation process is implemented through stochastic neural activity, directly aligning with the inherent noise in neuromorphic substrates [50, 45, 29] and the brain [11, 53, 37].
+Reinforcement learning mechanisms are realized through reward prediction errors (emulating dopaminergic reward signals [49, 48]), eligibility traces, and local synaptic plasticity.
+These elements also enable learning under uncertainty [17].
+This approach results in a learning rule that is both theoretically grounded and potentially realizable in event-driven neuromorphic hardware.
 Our results demonstrate the learning effectiveness of NRL, solving both immediate and delayed feedback/reward problems in simulated environments, achieving performance comparable to baselines optimized employing backpropagation.
-Additionally, it shows a significant improvement over RMHL [32, 40], a modern, well-known approach that employs noise as the core of its synaptic plasticity.
-Our scalability results further demonstrate that NRL effectively assigns credit even in multilayer networks, a challenge for many biologically plausible algorithms, indicating its potential for scaling.
-This is a critical difference with RMHL, which struggles to assign credit in multilayer networks evidenced by its minimal performance improvements across trials.
+In addition, it shows a significant improvement over RMHL [33, 38], a modern approach that also uses noise as the core of its synaptic plasticity.
+Our scalability results further demonstrate that NRL effectively assigns credit even in multilayer networks, a challenge for many biologically plausible or local algorithms, indicating its potential for scaling.
+This is a critical difference from RMHL, which struggles to assign credit in multilayer networks, as evidenced by its minimal performance improvements across trials.
 Nevertheless, as the depth of the network increases, NRL shows slower convergence due to its noise-driven updates.
 NRL shares similarities with node perturbation methods [18, 23, 65, 62, 15], which inject noise into neurons to adjust weights based on state and performance fluctuations.
-However, unlike node perturbation – which requires constant feedback or reward for learning – NRL is well-suited for delayed reward setups.
+However, unlike node perturbation, which requires constant feedback or reward for learning, NRL is well-suited for delayed reward setups.
 By leveraging eligibility traces [22, 25], it can effectively assign credit to past neural states even when actions and outcomes are separated in time.
 Gradient approximation, a cornerstone of our learning rule’s derivation, is also crucial in mathematical optimization and machine learning, enabling learning when exact gradients are computationally expensive or unavailable.
-Techniques such as forward gradient methods [2] and zeroth-order (ZO) optimization [35, 8] have emerged, utilizing stochastic perturbations and function value comparisons to estimate gradients.
+Techniques such as forward gradient methods [2] and zeroth-order optimization [34, 6] have emerged, utilizing stochastic perturbations and function value comparisons to estimate gradients.
 Recent advances, such as employing local auxiliary networks for informed gradient guesses [19], have significantly improved the alignment of these approximations with exact gradients, reducing the performance gap with backpropagation.
@@
 Furthermore, compared to traditional gradient-based methods like backpropagation, gradient approximation methods offer a key advantage in computationally constrained scenarios, or when using non-differentiable networks, such as spiking neural networks [66, 58].
-Additionally, the local-update structure of NRL provides compatibility with neuromorphic substrates [49], where efficient, low-power computations are essential.
-Neuromorphic systems can also inherently simulate noisy neural activity, enabling energy-efficient learning in applications where real-time adaptation is key.
-The inherent noise in NRL is also advantageous for reinforcement learning setups, as it naturally facilitates exploration-exploitation balance [27, 30].
-Injected noise perturbs neural states, promoting exploratory deviations from expected values and thus introducing variations in the agent’s policy.
-Our work has limitations that suggest future research directions.
-First, its dependence on multiple forward passes per synaptic update reduces biological plausibility, even though these passes could be performed in parallel.
+Additionally, the local-update structure of NRL provides direct compatibility with neuromorphic substrates [50], where only locally available information in the synapse is prioritized and global communication is severely penalized.
+This eliminates the need for the high-bandwidth global data movement and large memory buffers required by backpropagation, the primary source of energy consumption in deep learning models, addressing the fundamental energy and latency bottlenecks of conventional Von Neumann systems.
+Furthermore, NRL’s reliance on noise for gradient approximation offers a paradigm shift: it transforms the device stochasticity common in neuromorphic circuits from an engineering challenge into a computational resource.
+The inherent noise in NRL is also advantageous for reinforcement learning setups, as it naturally facilitates the exploration-exploitation balance [26, 31].
+Injected noise perturbs neural states, promoting exploratory deviations from expected values and thus introducing variations in the agent’s policy, which benefits adaptation on neuromorphic edge devices.
+Despite its promise, our work has limitations that suggest future research directions.
+The current requirement for multiple forward passes (noisy and clean) per synaptic update, while theoretically simple, introduces a time-latency overhead.
 Future work should explore more efficient ways to estimate noise impacts, potentially eliminating the need for multiple passes.
-Second, our modeling approach deliberately involves a high degree of abstraction: we use rate-based neurons instead of spiking models, do not distinguish between excitatory and inhibitory populations, and treat synaptic potentiation and depression as a single continuum.
-While this provides clear insight into the capabilities and limitations of our approach, more biologically detailed models could offer further insights.
-Third, our reward prediction error calculation is a simplification.
-While sufficient for demonstrating the learning mechanism, the simplicity of using a running average may limit its effectiveness in scenarios where navigating through transient low-reward regions is necessary to achieve higher long-term rewards.
-In such cases, the model may underestimate the value of actions that lead to delayed benefits, potentially resulting in suboptimal updates.
-More sophisticated approaches, such as temporal difference learning [52, 54], or learned prediction through a critic network, could provide a richer and more temporally sensitive estimate of expected rewards.
-These strategies, which align more closely with biological dopaminergic signaling and can be implemented within actor-critic architectures, represent promising directions for improving learning performance in future work.
+In addition, while we focused on a rate-based model, the deployment of NRL on physical hardware like Intel Loihi [8] requires its extension to spiking neural networks (SNNs).
+Furthermore, the simplicity of our RPE calculation limits its robustness in more complex tasks requiring long-term planning.
+Future development should incorporate more sophisticated temporal difference learning mechanisms [54, 56], such as a learned value function within an actor-critic architecture [28], to provide a richer temporally sensitive estimate of rewards.
 Finally, while backpropagation may offer faster convergence in unconstrained settings, recent studies [7, 13], including ours, indicate that this gap can be narrowed for specific networks and tasks.
-Our findings highlight the potential of unifying optimization methods and neurobiological principles.
-By drawing inspiration from both domains, we have developed a learning rule that not only offers a biologically plausible alternative to backpropagation but also provides insights into how learning, particularly synaptic plasticity, might occur in biological neural systems.
-Conversely, the constraints of biological realism can guide the development of more efficient and robust artificial learning algorithms, particularly suited for neuromorphic hardware and real-time applications.
-This synergistic approach promises to advance our understanding of intelligence in both biological and artificial systems [64, 38], driving progress in neuroscience and machine learning alike.
+Our findings ultimately highlight the potential for unifying optimization methods and neurobiological principles.
+By drawing inspiration from both domains, we have developed a learning rule that not only offers an alternative to backpropagation but is also algorithmically suited for low-energy, event-driven neuromorphic hardware.
+In the same way, the constraints of biological realism can guide the development of more efficient and robust artificial learning algorithms.
+This synergistic approach provides a crucial new path for creating low-energy, fast, and adaptive AI systems [68, 36].
 Acknowledgments
@@
 […]
-Now, substituting the expectation with a sampling under some empirical distribution,
+Now, substituting the expectation with a sampling under some empirical distribution:
 […]
-Finally, defining gradient descent on parameters $\theta$ as $\theta_{t+1}=\theta_{t}+\alpha\nabla g(\theta)$,
+Finally, defining gradient descent on parameters $\theta$ as $\theta_{t+1}=\theta_{t}+\alpha\nabla g(\theta)$:
 $\theta_{t+1}$ $=\theta_{t}+\alpha\nabla g(\theta)$

### v4

@@
 This contrasts sharply with the hardware needs for local synaptic updates that rely only on information available at the synaptic weight.
-In addition, BP’s iterative forward and backward passes are incompatible with the asynchronous event-driven regime in which most neuromorphic devices operate [41], and its deterministic nature cannot tolerate inherit noise in neuromorphic circuits.
+In addition, BP’s iterative forward and backward passes are incompatible with the asynchronous event-driven regime in which most neuromorphic devices operate [47], and its deterministic nature cannot tolerate inherent noise in neuromorphic circuits.
 In response, local plasticity approaches and biologically-inspired learning rules have been proposed to bypass the constraints of BP [73, 72, 35, 55, 25, 76].
 A particularly promising family of methods is noise-based learning, which leverages stochastic perturbations (either in weights [12, 78, 5, 9] or activations [19, 24, 74, 71, 15]) to approximate the true direction of the gradient.
-By injecting small random perturbations, these methods avoid the need for perfectly matched feedback pathways and can leverage inherit noise [51, 35].
+By injecting small random perturbations, these methods avoid the need for perfectly matched feedback pathways and can leverage inherent noise [60, 40].
 Given the prevalence of noise on neuromorphic and biological substrates [11, 62, 42], using noise as a mechanism to learn synaptic weights is an area of growing interest, with reward-modulated Hebbian learning (RMHL) [36, 43] as a promising candidate.
@@
 Despite theoretical advances, few noise-based methods have been adapted to real-world tasks.
-Some attempts which have been made include application to control problems [4], for spatiotemporal pattern generation [27], and for training non-differentiable spiking neural networks training [14], and these particularly struggle with delayed reward regimes.
+Some attempts which have been made include application to control problems [4], for spatiotemporal pattern generation [29], and for training non-differentiable spiking neural networks [14], and these particularly struggle with delayed reward regimes.
 This gap has limited the integration of noise-based approaches for effective credit assignment in temporally extended setups.
@@
 In practice, we consider $K=1$, analogous to single-sample updates in stochastic gradient descent.
+Although increasing $K$ (using multiple noise samples per step) would reduce the variance of the gradient estimate, it would linearly increase the computational cost and latency by requiring $K$ forward passes per update.
 Here, $g\bigl(\tilde{\theta}^{(i)}_{t}\bigr)=\log\pi_{\tilde{\theta}_{t}}(a_{t}\mid s_{t})$ and $g\left(\theta_{t}\right)=\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$ are the log of the noise-perturbed and noise-free output, respectively.
 Thus, we define $\rho_{t}=\log\pi_{\tilde{\theta}_{t}}(a_{t}\mid s_{t})-\log\pi_{\theta_{t}}(a_{t}\mid s_{t})$, which captures the impact of the noise on the policy.
+Computationally, this term measures the influence of the perturbation on the action selection, as it quantifies how much the injected noise increased or decreased the likelihood of the chosen action compared to the noise-free baseline.
 Putting this together for the $K=1$ case, we obtain our noise-based learning rule
@@
 Here, the general parameter vector $\theta_{t}=\{W^{1}_{t},\ldots,W^{L}_{t}\}$ corresponds to the full set of all weight matrices for all layers at time $t$.
-To do so, we first rewrite the gradient term as
-$\nabla g\left(W^{l}_{t}\right)=\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)$ (10)
-and apply the chain rule to break down the gradient with respect to $W^{l}_{t}$ as
+To do so, we first rewrite the gradient term as $\nabla g\left(W^{l}_{t}\right)=\nabla\log\pi_{\theta_{t}}\left(a_{t}\mid s_{t}\right)$ and apply the chain rule to break down the gradient with respect to $W^{l}_{t}$ as
 $\nabla g\left(W^{l}_{t}\right)=\nabla g\left(h^{l}_{t}\right)\frac{\partial h^{l}_{t}}{\partial W^{l}_{t}}=\nabla g\left(h^{l}_{t}\right)(x^{l-1}_{t})^{\top}.$ (10)
@@
 […]
-where $i$ indexes over a set of repeated samples of the noise term, $\xi$.
-Again, as above, by reducing this to the single sample ($n=1$) case, we can define a layer-specific weight update rule
+where $i$ indexes over a set of repeated samples of the noise term $\xi$. Again, as above, by reducing this to the single sample ($n=1$) case, we can define a layer-specific weight update rule
 $W^{l}_{t+1}\leftarrow W^{l}_{t}+\eta\delta_{t}\bar{\xi}^{l}_{t}{\rho}_{t}(\tilde{x}^{l-1}_{t})^{\top}$ (12)
@@
 In reality, however, rewards are often sparse, arriving only after a sequence of actions or upon reaching specific milestones.
-To handle this, we’ll modify our learning rule to account for rewards received at arbitrary times.
-Let us denote these discrete reward times as $\tau_{m}\in\{\tau_{0},\tau_{1},...\tau_{M}\}$. At these moments, we compute the RPE as
+To handle this, we will modify our learning rule to account for rewards received at arbitrary times.
+Let us denote these discrete reward times as $\tau_{m}\in\{\tau_{0},\tau_{1},\dots,\tau_{M}\}$. At these moments, we compute the RPE as
 $\delta_{\tau_{m}}=r_{\tau_{m}}-\bar{r}_{\tau_{m}}.$ (13)
@@
 We use implementations given by the libraries OpenAI Gym [3] and NeuroGym [45] for the different environments.
-We compare NRL to two baselines: an exact-gradient version of NRL, which serves as an ”optimal performance” benchmark, and an RMHL approach.
+We compare NRL to two baselines: an exact-gradient version of NRL, which serves as an “optimal performance” benchmark, and an RMHL approach.
 The first baseline, which we refer to as the exact-gradient method (BP baseline), is similar to an actor-only variant of actor-critic methods, relying on a running average of past rewards as the prediction error.
@@
 All updates are written in a per-layer form using $W^{l}$ and assume a delayed reward received at time $\tau_{m}$.
-Table 1: Learning rules. The terms in the BPTT update are rearranged for easier comparison.
+Table 1: Learning rules.
+The terms in the BPTT update are rearranged for easier comparison.
+Updates on NRL and RMHL rely on local synaptic information (inputs $\tilde{x}$, noise $\xi$) modulated by global broadcasted scalars (RPE $\delta$, and for NRL, noise impact $\rho$).
+In contrast, BPTT relies on non-local gradient information ($\nabla$) propagated from later layers.
+The updates are given by $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\Delta W^{l}_{\tau_{m}}$.
 Rule Update
-NRL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{\,l}_{t}\,\rho_{t}\,(\tilde{x}^{l-1}_{t})^{\top}$
-BPTT $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\delta_{\tau_{m}}\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})$
-RMHL $W^{l}_{\tau_{m}}\leftarrow W^{l}_{\tau_{m-1}}+\eta\;\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{\,l}_{t}\,(\tilde{x}^{l-1}_{t})^{\top}$
+NRL $\Delta W^{l}_{\tau_{m}}=\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\bar{\xi}^{\,l}_{t}\,\rho_{t}\,(\tilde{x}^{l-1}_{t})^{\top}$
+BPTT $\Delta W^{l}_{\tau_{m}}=\delta_{\tau_{m}}\nabla_{W^{l}_{\tau_{m-1}}}\log\pi_{\theta_{\tau_{m-1}}}(a_{t}\mid s_{t})$
+RMHL $\Delta W^{l}_{\tau_{m}}=\delta_{\tau_{m}}\!\sum_{t=\tau_{m-1}}^{\tau_{m}}\xi^{\,l}_{t}\,(\tilde{x}^{l-1}_{t})^{\top}$
 In environments with high reward variability, like Cartpole and Acrobot, we stabilize synaptic updates by scaling the RPE by dividing by $r_{\tau_{m}}$.
+This normalization accounts for the non-stationary nature of the return in some tasks, where the reward magnitude scales with trial duration.
+By treating the error as relative to the current return, we prevent update instability as the agent’s performance improves.
 Hyperparameter values and training details are provided in Appendix B.
@@
 Training details are provided in Appendix B.
-In Section titled “Scalability to deeper architectures”, we extend the comparison to deeper networks with multiple hidden layers.
+In Section “Scalability to deeper architectures”, we extend the comparison to deeper networks with multiple hidden layers.
 First, the Reaching problem [22], visualized in Fig.
@@
 To evaluate the accuracy of this approximation, we employ a single hidden-layer network and the Acrobot problem.
+We chose this task as it offers a representative delayed-reward challenge, unlike the Reaching task.
+Its stable dynamics also allowed us to clearly separate the approximation error from the variance inherent to the task itself, which was difficult with Cartpole’s sensitive initial conditions.
 The difference between the clean pass output and the averaged noisy passes output was calculated for each timestep and averaged over 500 timesteps, as shown in Fig.
@@
 Nevertheless, as the depth of the network increases, NRL shows slower convergence due to its noise-driven updates.
-NRL shares similarities with node perturbation methods [18, 23, 65, 62, 15], which inject noise into neurons to adjust weights based on state and performance fluctuations.
-However, unlike node perturbation, which requires constant feedback or reward for learning, NRL is well-suited for delayed reward setups.
-By leveraging eligibility traces [22, 25], it can effectively assign credit to past neural states even when actions and outcomes are separated in time.
+Beyond network depth, NRL performance is also influenced by the duration of the reward delay.
+As this delay increases, the eligibility trace accumulates more noise, lowering the signal-to-noise ratio in the gradient estimate.
+Consequently, although our RPE normalization prevents instability caused by large reward values in long episodes, we expect convergence to slow down as the delay extends, reflecting the difficulty of assigning credit over long time windows.
+This challenge can be compared with the vanishing gradient problem in BPTT, characterized by a signal decay.
+Yet, it contrasts with the additive variance accumulation problem seen in NRL, where the learning signal is not lost, but rather obscured by the noise of non-causal perturbations.
+Regarding the scope of applicability, our framework suggests specific behaviors in complex scenarios.
+For high-dimensional observation spaces, the main challenge for perturbation-based learning is the potential variance increase in the gradient estimate.
+In this scenario, NRL may require more samples for convergence.
+However, the inclusion of a baseline via the RPE serves as a variance reduction mechanism [74], stabilizing the learning signal even in larger input spaces.
+When it comes to handling continuous action spaces, NRL, as a policy gradient approximation, is naturally compatible with continuous control.
+This extension would not require any changes in the learning rule, but simply replacing the categorical output with a Gaussian policy head, where noise perturbs the action mean directly.
+NRL shares similarities with node perturbation methods [19, 24, 74, 71, 15], which also use noise to guide learning.
+However, they differ in their optimization objective and temporal scope.
+Standard node perturbation is formulated as a zeroth-order optimization technique for estimating the gradient of an immediate loss function.
+In contrast, NRL is derived as an approximation of policy gradients, designed to maximize long-term expected return.
+This distinction is realized through the use of an eligibility trace [23, 26], which keeps a temporary memory of local information (perturbation, activations, etc) until a delayed reward arrives.
+This effectively assigns credit to past neural states when actions and outcomes are separated in time.
+In addition, NRL also incorporates variance reduction via the RPE ($\delta_{t}=r_{t}-\bar{r}_{t}$), where the expected reward $\bar{r}_{t}$ acts as a baseline to stabilize the gradient estimate.
 Gradient approximation, a cornerstone of our learning rule’s derivation, is also crucial in mathematical optimization and machine learning, enabling learning when exact gradients are computationally expensive or unavailable.
@@
 Furthermore, NRL’s reliance on noise for gradient approximation offers a paradigm shift: it transforms the device stochasticity common in neuromorphic circuits from an engineering challenge into a computational resource.
-The inherent noise in NRL is also advantageous for reinforcement learning setups, as it naturally facilitates the exploration-exploitation balance [26, 31].
-Injected noise perturbs neural states, promoting exploratory deviations from expected values and thus introducing variations in the agent’s policy, which benefits adaptation on neuromorphic edge devices.
+The inherent noise in NRL is also advantageous for reinforcement learning setups, as it naturally introduces the stochasticity required for exploration [28, 34].
+Injected noise perturbs neural states, promoting deviations from expected values and thus introducing variations in the agent’s policy.
+Recent studies on noise-based adaptation [12] have demonstrated that noise-driven learning mechanisms can effectively balance exploration and exploitation, particularly in volatile environments.
 Despite its promise, our work has limitations that suggest future research directions.
-The current requirement for multiple forward passes (noisy and clean) per synaptic update, while theoretically simple, introduces a time-latency overhead.
-Future work should explore more efficient ways to estimate noise impacts, potentially eliminating the need for multiple passes.
+The current requirement for both noisy and clean forward passes (or averaging multiple noisy passes) per synaptic update, while theoretically simple, introduces a time-latency and computational overhead.
+In practice, this overhead is mitigated by the accelerated dynamics of neuromorphic hardware, which typically operates at microsecond timescales, orders of magnitude faster than real-world environments [8].
+Thus, performing multiple forward passes potentially has a negligible impact on latency.
+From an energy perspective, the cost of extra passes is offset by eliminating the high-bandwidth data transport and memory buffering required by gradient-based methods, such as BP [13], which remain the primary energy bottlenecks in conventional learning systems.
+However, to further enhance this method, future work could explore alternative ways to estimate noise impacts, eliminating the need for multiple passes.
 In addition, while we focused on a rate-based model, the deployment of NRL on physical hardware like Intel Loihi [8] requires its extension to spiking neural networks (SNNs).
+Adapting NRL to this domain presents specific implementation challenges.
+For instance, the continuous eligibility traces used here would need to be replaced by synaptic tracers that accumulate discrete spike events via exponential decay kernels [26].
+Similarly, the injected noise used for gradient approximation could translate into stochastic membrane potential fluctuations or probabilistic firing thresholds [40], transforming perturbations into spike-timing variability.
+Also, while our derivation assumes Gaussian noise, physical hardware often exhibits non-Gaussian stochasticity.
+However, as elaborated in Appendix C, the central limit theorem suggests that the aggregation of these discrete noise sources at the neuronal level effectively approaches a Gaussian distribution.
+Thus, our theoretical framework is expected to remain robust in physical neuromorphic systems.
 Furthermore, the simplicity of our RPE calculation limits its robustness in more complex tasks requiring long-term planning.
@@
 Finally, while backpropagation may offer faster convergence in unconstrained settings, recent studies [7, 13], including ours, indicate that this gap can be narrowed for specific networks and tasks.
+Finally, regarding sample efficiency, NRL is an on-policy algorithm theoretically grounded in the policy gradient theorem [66].
+Although off-policy methods can improve data efficiency, they typically rely on mechanisms that conflict with biological and neuromorphic constraints.
+Incorporating off-policy mechanisms, such as experience replay [44] or importance sampling [52], would require centralized memory buffers and global data transport, violating the principle of synaptic locality.
+Crucially, this restriction allows us to isolate and evaluate the performance of the learning rule itself, avoiding the obscuring factors introduced by add-ons.
+Future work could investigate how such off-policy augmentations can be adapted to the neuromorphic context to further enhance sample efficiency.
 Our findings ultimately highlight the potential for unifying optimization methods and neurobiological principles.
@@
 The number of episodes was chosen to illustrate the convergence of our method relative to baseline methods, with 8000, 20000, and 1000 episodes used for the Acrobot, Cartpole, and Reaching problems, respectively.
-In all our experiments, we set the smoothing factor for reward estimation to $\lambda=0.66$, striking a balance between recent values (for quick adaptability) and a longer history (for robustness against rapid reward fluctuations).
+In all our experiments, we set the smoothing factor for reward estimation to $\lambda=0.66$.
+This value was empirically found to offer a robust trade-off across all tasks, striking a balance between recent values (for quick adaptability) and a longer history (for robustness against rapid reward fluctuations).
+We observed that this choice was sufficiently efficient to be applied uniformly across all environments without the need for task-specific adjustment.
 Each neural network consisted of an input, hidden, and output layer.
@@
 $\sigma$ - 1e-3 1e-1
+Appendix C Noise distributions in neuromorphic hardware
+In our derivation of the NRL update rule, we assume that the perturbative noise $\xi$ follows a Gaussian distribution.
+This assumption is crucial for maintaining equality between the expected update and the true gradient.
+However, in physical neuromorphic implementations, noise sources at the device level may follow different distributions.
+In practice, neuromorphic substrates manifest stochasticity through physical processes.
+Thermal or Johnson–Nyquist noise [27], arising from charge carriers, is inherently Gaussian and is always present in analog circuits.
+However, other sources are non-Gaussian.
+Shot noise [56], caused by the discrete nature of electric charge crossing a barrier, follows a Poisson distribution.
+Random telegraph noise [30], common in memristive devices, manifests as discrete switching between two states, following a bimodal distribution.
+Although individual synaptic or neuronal noise sources may follow non-Gaussian distributions, the NRL update rule operates on the aggregated activity influencing the neuron’s state.
+A fundamental property of neural integration is the summation of inputs from multiple presynaptic sources.
+According to the central limit theorem (CLT) [17], the sum of a large number of independent random variables, regardless of their original distribution, provided they have finite variance, converges to a Gaussian distribution.
+Consequently, the Gaussian assumption in $\xi$ used in our derivation is not only a mathematical convenience, but a statistically valid approximation of the effective aggregate noise seen in hardware implementations or biological networks.
+Even in scenarios with low connectivity where the CLT approximation is not perfect, the learning rule is expected to remain robust.
+While Theorem 25 relies on this Gaussian assumption for exact equality between the expected update and the true gradient, in practice, optimization does not require the update vector to be identical to the true gradient, but to have a positive projection onto it, i.e., a positive cosine similarity.
+This principle is well-established in the literature, where algorithms, such as feedback alignment [38], direct feedback alignment [48], or sign-symmetry [37], demonstrate that approximate gradients are sufficient for convergence, as long as they remain within $90^{\circ}$ of the steepest descent direction.
+Therefore, skewed non-Gaussian noise may reduce convergence speed by lowering this cosine similarity, but it does not orthogonalize the update vector, thus maintaining a valid descent trajectory.